### v1

# Gauge-dependence of Scalar Induced Gravitational Waves

Abstract
In this review we look into the gauge-dependence of scalar-induced gravitational waves (SIGWs) that are second-order tensors produced by first-order scalar-modes. The method includes deriving the background, first- and second-order Einstein field equations without imposing a gauge. We address the gauge-invariant approach and study the source-term of SIGWs in three different gauges, synchronous, Poisson and uniform curvature gauge. We find that numerically computed kernels in all three gauges behave closely with minimal discrepancy. As expected, when going in sub-horizon modes, $x\gg 1$, the discrepancy decreases and the behavior matches, pointing to a gauge-invariant observable.
keywords:
gravitational waves / theory, gauge, primordial gravitational waves, second-order perturbation theory
1 Introduction
Due to the weak interaction with matter, we anticipate to detect primordial gravitational waves (GWs) with current- and future-generation detectors [1]. Sources of GWs from the early universe have been extensively studied in literature [2, 3, 4] and of the various sources one contribution arises from “scalar-induced” GWs (SIGWs) [5, 6, 7, 8, 9, 10]. These are observed as a second-order source of GWs produced by the coupling of large scalar fluctuations at horizon re-entry [11, 12]. The frequency range of SIGWs produced in the radiation dominated era fall within current and future generation ground- and space-based detectors, for example Laser Interferometer Space Antenna (LISA) [13, 14] and Pulsar Timing Array (PTA) [15, 16].
A perturbative approach is used to study SIGWs, where small inhomogeneous perturbations on a homogeneous and isotropic Friedmann-Roberston-Walker (FRW) background are considered [17]. This splitting of an unperturbed background and perturbed system introduces the complication in gauge choice. There is no preferred choice of coordinate in relativistic perturbation theory, allowing one to work in the appropriate gauge best suited for the problem. It is well known that at linear-order, scalar modes are gauge-dependent while tensor modes are gauge-independent, i.e. do not change under coordinate transformation [18]. However due to the non-linear feature of the source, the second-order tensor modes that are sourced by these first-order scalar modes must also depend on the gauge conditions [19]. In previous studies of “scalar-induced” tensor modes, a gauge condition is usually imposed beforehand, usually the zero-shear gauge, also knows as longitudinal or Poisson gauge.
Detailed calculations has been done in previous studies presenting the gauge-dependence of induced GWs by studying different gauges [20, 21, 22]. There have been various ways in approaching this issue. One approach has been by constructing a gauge-invariant formalism for the second-order tensor modes by the addition of terms with the intention to reduce the gauge dependence [23, 24, 25, 26]. However there are some complications with this approach, as pointed out by [27, 28], regarding the understanding between the observable and the gauge-invariant variable. They mention that there is not much difference between choosing particular gauge from the beginning and working with specific gauge-invariant combination.
Instead, the second approach in tackling the gauge issue is by defining a gauge suitable to best describe the observable [27, 29, 28]. In [27], they argue that the suitable gauge is one in which the coordinates are fixed according to the positions of the interferometer mirrors. This is known as the synchronous gauge, or TT gauge. [27] show that the observable GW spectrum of GWs in the TT gauge and Poisson gauge coincide. They discuss that this is a result of the GWs generated during the radiation-dominated era being treated as linear perturbations of the metric as they are freely propagating inside the horizon. Therefore the gauge-dependence does not remain. A similar argument was made by [28], in which they show that the scalar-induced gravitational waves behave similar when calculated in gauges which are “well-behaved” on small scales. The authors point out that the Newtonian gauge is suitable both for the calculation and the physical interpretation. Further mentioning other gauges, among them the synchronous gauge, to also be “well-behaved” in small scales in a universe dominated by perfect fluid with constant equations of state.
In this work, we review the derivation of the source term of scalar-induced gravitational waves in a generic gauge and further on revisit the gauge issue. In Sections 2 and 3, we introduce the general framework for cosmological perturbations of the metric and then derive the Einstein field equations. In Section 3.3, the source term is presented without imposing a gauge. Further on the gauge transformation and gauge-invariant variables are presented in Section 4. We work with the approaches mentioned to show the behaviour of the scalar-induced gravitational waves in different gauges. In Sections 5 and 6, we study the source-term of the SIGWs and the subsequent kernel in three different gauges.
In the following calculations, we will be using natural units, $c=\hbar=1$ and Planck mass $M_{p}^{-2}=8\pi G=\kappa^{2}$. In addition, we are considering conformal time, $\tau$, where $a(\tau)d\tau=dt$. To denote four-dimensional spacetime indices, we have used Greek indices, $\mu,\nu$. While Roman indices are used to denote spatial indices $i,j$. The total space-time metric has signature (,+,+,+)(-,+,+,+)( - , + , + , + ).
2 Cosmological perturbation of the metric
Any tensorial quantity, $\mathbf{A}$, can be decomposed into background and inhomogeneous parts,
$\mathbf{A}=\mathbf{A}^{(0)}(\tau)+\delta^{(r)}\mathbf{A}(\tau,x^{i}).$ (1)
The background quantity is $\mathbf{A}^{(0)}$ and only has time-dependence. The perturbed quantity up to $r$-th order, $\delta^{(r)}\mathbf{A}$, is defined as
$\delta^{(r)}\mathbf{A}=\sum_{r=1}^{\infty}\frac{1}{r!}\delta^{(r)}\mathbf{A},$ (2)
where it has both time- and spatial-dependence. The metric tensor, $g_{\mu\nu}$, up to the second order is
$g_{\mu\nu}=g_{\mu\nu}^{(0)}+\delta^{(1)}g_{\mu\nu}+\frac{1}{2}\delta^{(2)}g_{\mu\nu},$ (3)
where $\delta g_{\mu\nu}$ and $\delta^{(2)}g_{\mu\nu}$ is the perturbed metric in first and second order. Using a method known as (3+1) decomposition, 3-dimensional quantities are split into scalar, vector and tensor components corresponding to the transformational behavior on spatial hypersurfaces. The components of the perturbed spatially flat Robertson - Walker metric can be written as
$g_{00}$ $=-a^{2}(\tau)\left(1+2\sum_{r=1}^{\infty}\frac{1}{r!}\phi^{(r)}(\vec{x},\tau)\right),$
$g_{0i}$ $=a^{2}(\tau)\sum_{r=1}^{\infty}\frac{1}{r!}\hat{w}_{i}^{(r)}(\vec{x},\tau),$
$g_{ij}$ $=a^{2}(\tau)\left\{\left[1-2\left(\sum_{r=1}^{\infty}\frac{1}{r!}\psi^{(r)}(\vec{x},\tau)\right)\right]\delta_{ij}+\sum_{r=1}^{\infty}\frac{1}{r!}\hat{h}_{ij}^{(r)}(\vec{x},\tau)\right\},$ (4)
where the $0-i$ and $i-j$ components can be further decomposed,
$\hat{w}_{i}^{(r)}=\partial_{i}w^{(r)}+w_{i}^{(r)},$ (5)
$\hat{h}_{ij}^{(r)}=D_{ij}h^{(r)}+\partial_{i}h_{j}^{(r)}+\partial_{j}h_{i}^{(r)}+h_{ij}^{(r)}.$ (6)
The scalar components are thus $\psi,\phi$, $w$ and $h$, the vectors $w_{i}$ and $h_{i}$ are transverse, i.e. $\partial^{i}w_{i}^{(r)}=\partial^{i}h_{i}^{(r)}=0$, and $h_{ij}^{(r)}$ is a symmetric transverse and trace-free tensor, i.e. $\partial^{i}h_{ij}^{(r)}=h_{i}^{i(r)}=0$. Finally, $D_{ij}=\partial_{i}\partial_{j}-\frac{1}{3}\nabla^{2}\delta_{ij}$ is the trace-free operator.
In order to study second-order tensor modes, the metric Section 2 can be simplified further with the following assumptions; (1) first-order vector perturbations have decreasing amplitude, thus being able to neglect $w^{(1)}_{i}$ and $h^{(1)}_{i}$, (2) first-order tensors give negligible contribution to second-order perturbations, hence neglecting $h_{ij}^{(1)}$. On the other hand, we will keep the second-order vector and tensor perturbations since they can be generated by first-order scalar perturbations, i.e. the tensor modes being “scalar-induced” gravitational waves. With these assumptions the metric, Section 2, can be re-written as
$g_{00}$ $=-a^{2}(\tau)\left(1+2\phi^{(1)}+\phi^{(2)}\right),$
$g_{0i}$ $=a^{2}(\tau)\left(\partial_{i}w^{(1)}+\frac{1}{2}\partial_{i}w^{(2)}+\frac{1}{2}w_{i}^{(2)}\right),$
$g_{ij}$ $=a^{2}(\tau)\left[\left(1-2\psi^{(1)}-\psi^{(2)}\right)\delta_{ij}+D_{ij}\left(h^{(1)}+\frac{1}{2}h^{(2)}\right)\right.$
$\left.+\frac{1}{2}\left(\partial_{i}h_{j}^{(2)}+\partial_{j}h_{i}^{(2)}+h_{ij}^{(2)}\right)\right].$ (7)
The contra-variant metric tensor can be obtained using the following relation, $g_{\mu\nu}g^{\nu\lambda}=\delta^{\lambda}_{\mu}$, where
$g^{00}$ $=-a^{2}(\tau)\left(1-2\phi^{(1)}-\phi^{(2)}+4\left(\phi^{(1)}\right)^{2}-\partial^{i}w^{(1)}\partial_{i}w^{(1)}\right),$
$g^{0i}$ $=a^{2}(\tau)\left[\partial^{i}w^{(1)}+\frac{1}{2}\left(\partial^{i}w^{(2)}+w^{i(2)}\right)+2\left(\psi^{(1)}-\phi^{(1)}\right)\partial^{i}w^{(1)}\right.$
$\left.-\partial^{k}w^{(1)}D^{i}_{k}h^{(1)}\right],$
$g^{ij}$ $=a^{2}(\tau)\left[\left(1+2\psi^{(1)}+\psi^{(2)}+4\left(\psi^{(1)}\right)^{2}\right)\delta^{ij}+D^{ij}\left(h^{(1)}+\frac{1}{2}h^{(2)}\right)\right.$
$\left.-\frac{1}{2}\left(\partial^{i}h^{j(2)}+\partial^{j}h^{i(2)}+h^{ij(2)}\right)-\partial^{i}w^{(1)}\partial^{j}w^{(1)}\right.$
$\left.-4\psi^{(1)}D^{ij}h^{(1)}+D^{ik}h^{(1)}D^{j}_{k}h^{(1)}\right].$ (8)
The entire perturbed metric line-element is
$ds^{2}=a^{2}(\tau)$ $\left[-\left(1+2\phi^{(1)}+\phi^{(2)}\right)d\tau^{2}+\left(\partial_{i}w^{(1)}+\frac{1}{2}\partial_{i}w^{(2)}+\frac{1}{2}w_{i}^{(2)}\right)d\tau dx^{i}\right.$
$\left.\left\{\left(1-2\psi^{(1)}-\psi^{(2)}\right)\delta_{ij}+D_{ij}\left(h^{(1)}+\frac{1}{2}h^{(2)}\right)\right.\right.$
$\left.\left.+\frac{1}{2}\left(\partial_{i}h_{j}^{(2)}+\partial_{j}h_{i}^{(2)}+h_{ij}^{(2)}\right)\right\}dx^{i}dx^{j}\right].$ (9)
3 Einstein field equations
The Einstein field equation is defined as
$G^{\mu}_{\nu}=\kappa^{2}T^{\mu}_{\nu}.$ (10)
In the following sub-sections we study, background, first- and second-order Einsteins equations in a generic gauge. The components in order to find the field equations is presented in appendix A, B, C, D.
3.1 Background Einsteins equations
Using the background zeroth-order Einstein equation, the following terms are defined
$\mathcal{H}^{2}$ $=\frac{\kappa^{2}a^{2}}{3}\bar{\rho},$
$\mathcal{H}^{\prime}$ $=-\frac{\kappa^{2}a^{2}}{6}\left(\bar{\rho}+3\bar{P}\right),$ (11)
where $\bar{P}$ and $\bar{\rho}$ are the background quantities for pressure and energy-density and $\mathcal{H}=\frac{a^{\prime}}{a}$ is the conformal Hubble parameter.
3.2 First order Einsteins equations
The first-order Einstein field equations are used to find expressions for first-order perturbed matter quantities. Using the time-time component, the expression for first-order energy density is found
$G^{(1)0}_{0}$ $=\kappa^{2}T^{(1)0}_{0},$
$a^{-2}\left[6\mathcal{H}^{2}\phi^{(1)}+6\mathcal{H}\psi^{(1)^{\prime}}+2\mathcal{H}^{2}\nabla^{2}w^{(1)}-2\nabla^{2}\psi^{(1)}-\frac{1}{2}\partial_{k}\partial^{i}D^{k}_{i}h^{(1)}\right]$ $=-\kappa^{2}\rho^{(1)},$ (12)
$\rho^{(1)}=-\frac{1}{\kappa^{2}a^{2}}\left[6\mathcal{H}\left(\mathcal{H}\phi^{(1)}+\psi^{(1)^{\prime}}\right)+2\mathcal{H}^{2}\nabla^{2}w^{(1)}-2\nabla^{2}\psi^{(1)}-\frac{1}{2}\partial_{k}\partial^{i}D^{k}_{i}h^{(1)}\right].$ (13)
The time-space component provides the first-order 4-velocity expression
$G^{(1)0}_{i}$ $=\kappa^{2}T^{(1)0}_{i},$
$a^{-2}\left(-2\mathcal{H}\partial_{i}\phi^{(1)}-2\partial_{i}\psi^{(1)^{\prime}}-\frac{1}{2}\partial_{k}D^{k}_{i}h^{(1)^{\prime}}\right)$ $=\kappa^{2}\left(\bar{\rho}+\bar{P}\right)\left(\partial_{i}w^{(1)}+v^{(1)}_{i}\right),$ (14)
$v_{i}^{(1)}=-\frac{2}{\kappa^{2}a^{2}\left(\bar{\rho}+\bar{P}\right)}\left(\mathcal{H}\partial_{i}\phi^{(1)}+\partial_{i}\psi^{(1)^{\prime}}+\frac{1}{4}\partial_{k}D^{k}_{i}h^{(1)^{\prime}}\right)-\partial_{i}w^{(1)}.$ (15)
We can find the trace and trace-free part of spatial component, $G^{(1)i}_{j}=\kappa^{2}T^{(1)i}_{j}$ ,
$-a^{-2}$ $\left(2\mathcal{H}\phi^{(1)^{\prime}}+4\left[\mathcal{H}^{\prime}+\mathcal{H}^{2}\right]\phi^{(1)}-2\mathcal{H}^{2}\phi^{(1)}+4\mathcal{H}\psi^{(1)^{\prime}}\right.$
$\left.+2\psi^{(1)^{\prime\prime}}+2\nabla^{2}\left(\phi^{(1)}-\psi^{(1)}+2w^{(1)}+w^{(1)^{\prime}}\right)\right)=\kappa^{2}P^{(1)},$ (16)
$-a^{-2}\left(\partial^{i}\partial_{j}-\frac{1}{3}\nabla^{2}\delta^{i}_{j}\right)\left(\phi^{(1)}-\psi^{(1)}+2\mathcal{H}w^{(1)}+w^{(1)^{\prime}}\right)=\kappa^{2}a^{-2}\pi^{(1)i}_{j},$ (17)
where we can thus define the anisotropic stress as,
$\pi^{(1)i}_{j}=-\frac{1}{\kappa^{2}}\left(\partial^{i}\partial_{j}-\frac{1}{3}\nabla^{2}\delta^{i}_{j}\right)\left(\phi^{(1)}-\psi^{(1)}+2\mathcal{H}w^{(1)}+w^{(1)^{\prime}}\right).$ (18)
3.3 Second-order Einstein field equation and source term
For understanding the source-term of scalar-induced gravitational waves, we will be studying the spatial component of second-order Einstein tensor, specifically the transverse, trace-free part. We do this by first applying the projection tensor $P^{li}_{jm}$ to the field equation [30], i.e. $P^{li}_{jm}G^{(2)m}_{l}=\kappa^{2}P^{li}_{jm}T^{(2)m}_{l}$. Accordingly, the second-order vector and tensor modes are removed. Hence the spatial Einstein tensor, Section C.3, and energy-momentum tensor, Section D.3, can be written as
$P^{li}_{jm}G^{(2)m}_{l}$ $=a^{-2}\left[\frac{1}{4}h^{i(2)^{\prime\prime}}_{j}+\frac{1}{2}\frac{a^{\prime}}{a}h^{i(2)^{\prime}}_{j}-\frac{1}{4}\nabla^{2}h^{i(2)}_{j}+\partial^{i}\phi^{(1)}\partial_{j}\phi^{(1)}+2\phi^{(1)}\partial^{i}\partial_{j}\phi^{(1)}\right.$
$\left.-2\psi^{(1)}\partial^{i}\partial_{j}\phi^{(1)}-\partial_{j}\phi^{(1)}\partial^{i}\psi^{(1)}-\partial^{i}\phi^{(1)}\partial_{j}\psi^{(1)}+3\partial^{i}\psi^{(1)}\partial_{j}\psi^{(1)}\right.$
$\left.+4\psi^{(1)}\partial^{i}\partial_{j}\psi^{(1)}+2\frac{a^{\prime}}{a}\partial^{i}w^{(1)}\partial_{j}{\phi^{(1)}}+4\frac{a^{\prime}}{a}\phi^{(1)}\partial^{j}\partial_{i}w^{(1)}+\phi^{(1)^{\prime}}\partial^{i}\partial_{j}w^{(1)}\right.$
$\left.+2\phi^{(1)}\partial^{i}\partial_{j}w^{(1)^{\prime}}+\nabla^{2}w^{(1)}\partial^{i}\partial_{j}w^{(1)}-\partial_{j}\partial^{k}w^{(1)}\partial^{i}\partial_{k}w^{(1)}\right.$
$\left.-2\frac{a^{\prime}}{a}\partial^{i}\psi^{(1)}\partial_{j}w^{(1)}-2\frac{a^{\prime}}{a}\partial^{i}w^{(1)}\partial_{j}\psi^{(1)}-\partial^{i}\psi^{(1)^{\prime}}\partial_{j}w^{(1)}\right.$
$\left.+\partial^{i}w^{(1)}\partial_{j}\psi^{(1)^{\prime}}-\partial^{i}\psi^{(1)}\partial_{j}w^{(1)^{\prime}}-\partial^{i}w^{(1)^{\prime}}\partial_{j}\psi^{(1)}\right.$
$\left.-2\psi^{(1)}\partial^{i}\partial_{j}w^{(1)^{\prime}}+\psi^{(1)^{\prime}}\partial^{i}\partial_{j}w^{(1)}-4\frac{a^{\prime}}{a}\psi^{(1)}\partial^{i}\partial_{j}w^{(1)}\right.$
$\left.-2\frac{a^{\prime}}{a}\phi^{(1)}D^{i}_{j}h^{(1)^{\prime}}-\frac{1}{2}\phi^{(1)^{\prime}}D^{i}_{j}h^{(1)^{\prime}}-\phi^{(1)}D^{i}_{j}h^{(1)^{\prime\prime}}\right.$
$\left.+\frac{1}{2}\partial_{k}\phi^{(1)}\partial^{i}D^{k}_{j}h^{(1)}+\frac{1}{2}\partial_{k}\phi^{(1)}\partial_{j}D^{ki}h^{(1)}-\frac{1}{2}\partial_{k}\phi^{(1)}\partial^{k}D^{i}_{j}h^{(1)}\right.$
$\left.+\partial_{j}\partial_{k}\phi^{(1)}D^{ki}h^{(1)}+\frac{1}{2}\psi^{(1)^{\prime}}D^{i}_{j}h^{(1)^{\prime}}+\psi^{(1)^{\prime\prime}}D^{i}_{j}h^{(1)}\right.$
$\left.+2\frac{a^{\prime}}{a}\psi^{(1)^{\prime}}D^{i}_{j}h^{(1)}+\frac{1}{2}\partial_{k}\psi^{(1)}\partial^{i}D^{k}_{j}h^{(1)}+2\frac{a^{\prime}}{a}\psi^{(1)}D^{i}_{j}h^{(1)^{\prime}}\right.$
$\left.+\psi^{(1)}D^{i}_{j}h^{(1)^{\prime\prime}}+\frac{1}{2}\partial_{k}\psi^{(1)}\partial_{j}D^{ki}h^{(1)}-\frac{3}{2}\partial_{k}\psi^{(1)}\partial^{k}D^{i}_{j}h^{(1)}\right.$
$\left.+2\psi^{(1)}\partial_{k}\partial^{i}D^{k}_{j}h^{(1)}+2\psi^{(1)}\partial_{k}\partial_{j}D^{ki}h^{(1)}-2\psi^{(1)}\nabla^{2}D^{i}_{j}h^{(1)}\right.$
$\left.-\nabla^{2}\psi^{(1)}D^{i}_{j}h^{(1)}+\partial^{i}\psi^{(1)}\partial_{k}D^{k}_{j}h^{(1)}+\partial_{j}\psi^{(1)}\partial_{k}D^{ki}h^{(1)}\right.$
$\left.+\partial_{k}\partial^{i}\psi^{(1)}D^{k}_{j}h^{(1)}+\frac{1}{2}\partial^{i}w^{(1)}\partial_{k}D^{k}_{j}h^{(1)^{\prime}}+\frac{1}{2}\partial_{k}\partial^{i}w^{(1)}D^{k}_{j}h^{(1)^{\prime}}\right.$
$\left.+\frac{1}{2}\partial_{k}\partial_{j}w^{(1)}D^{ki}h^{(1)^{\prime}}-\frac{1}{2}\nabla^{2}w^{(1)}D^{i}_{j}h^{(1)^{\prime}}+\frac{1}{2}\partial^{k}w^{(1)}\partial^{i}D_{kj}h^{(1)^{\prime}}\right.$
$\left.+\frac{1}{2}\partial^{k}w^{(1)}\partial_{j}D^{i}_{k}h^{(1)^{\prime}}-\partial^{k}w^{(1)}\partial_{k}D^{i}_{j}h^{(1)^{\prime}}+\frac{1}{2}\partial^{k}w^{(1)^{\prime}}\partial^{i}D_{kj}h^{(1)}\right.$
$\left.+\frac{1}{2}\partial^{k}w^{(1)^{\prime}}\partial_{j}D^{i}_{k}h^{(1)}-\frac{1}{2}\partial^{k}w^{(1)^{\prime}}\partial_{k}D^{i}_{j}h^{(1)}+\partial_{k}\partial_{j}w^{(1)^{\prime}}D^{ik}h^{(1)}\right.$
$\left.+\frac{a^{\prime}}{a}\partial^{k}w^{(1)}\partial^{i}D_{kj}h^{(1)}+\frac{a^{\prime}}{a}\partial^{k}w^{(1)}\partial_{j}D^{i}_{k}h^{(1)}-\frac{a^{\prime}}{a}\partial^{k}w^{(1)}\partial_{k}D^{i}_{j}h^{(1)}\right.$
$\left.+2\frac{a^{\prime}}{a}\partial_{k}\partial_{j}w^{(1)}D^{ik}h^{(1)}-\frac{1}{2}D^{ki}h^{(1)^{\prime}}D_{kj}h^{(1)^{\prime}}-\frac{1}{2}\partial^{i}D_{mj}h^{(1)}\partial_{k}D^{km}h^{(1)}\right.$
$\left.-\frac{1}{2}\partial_{j}D^{i}_{m}h^{(1)}\partial_{k}D^{km}h^{(1)}+\frac{1}{2}\partial_{m}D^{i}_{j}h^{(1)}\partial_{k}D^{km}h^{(1)}\right.$
$\left.-\frac{1}{2}\partial_{k}\partial_{j}D^{i}_{m}h^{(1)}D^{km}h^{(1)}+\frac{1}{2}\partial_{k}\partial_{m}D^{i}_{j}h^{(1)}D^{km}h^{(1)}\right.$
$\left.+\frac{1}{2}D^{km}h^{(1)}\partial^{i}\partial_{j}D_{km}h^{(1)}-\frac{1}{2}\partial_{k}\partial^{i}D_{mj}h^{(1)}D^{km}h^{(1)}\right.$
$\left.+\frac{1}{4}\partial^{i}D^{mk}h^{(1)}\partial_{j}D_{mk}h^{(1)}-\frac{a^{\prime}}{a}D_{kj}h^{(1)^{\prime}}D^{ik}h^{(1)}-\frac{1}{2}D_{kj}h^{(1)^{\prime\prime}}D^{ki}h^{(1)}\right.$
$\left.-\partial_{m}\partial_{k}D^{m}_{j}h^{(1)}D^{ki}h^{(1)}+\frac{1}{2}\partial_{m}\partial^{m}D_{kj}h^{(1)}D^{ki}h^{(1)}\right.$
$\left.+\frac{1}{2}\partial_{m}D^{ik}h^{(1)}\partial^{m}D_{kj}h^{(1)}-\frac{1}{2}\partial_{m}D^{ik}h^{(1)}\partial_{k}D^{m}_{j}h^{(1)}\right],$ (19)
$P^{li}_{jm}T^{(2)m}_{l}=$ $\frac{4}{a^{4}\kappa^{4}\left(\bar{\rho}+\bar{P}\right)}\left[\partial^{i}\left(\psi^{(1)^{\prime}}+\mathcal{H}\phi^{(1)}\right)\partial_{j}\left(\psi^{(1)^{\prime}}+\mathcal{H}\phi^{(1)}\right)\right]$
$+\frac{1}{a^{4}\kappa^{4}\left(\bar{\rho}+\bar{P}\right)}\left[\partial^{i}\left(\psi^{(1)^{\prime}}+\mathcal{H}\phi^{(1)}\right)\partial_{k}D^{k}_{j}h^{(1)^{\prime}}\right.$
$\left.+\partial_{j}\left(\psi^{(1)^{\prime}}+\mathcal{H}\phi^{(1)}\right)\partial_{k}D^{ik}h^{(1)^{\prime}}\right]$
$+\frac{1}{2a^{4}\kappa^{4}\left(\bar{\rho}+\bar{P}\right)}\partial_{k}D^{ik}h^{(1)^{\prime}}\partial_{k}D^{k}_{j}h^{(1)^{\prime}}$
$+\frac{2}{a^{2}\kappa^{2}}\left[\partial^{i}w^{(1)}\mathcal{H}\partial_{j}\phi^{(1)}+\partial^{i}w^{(1)}\partial_{j}\psi^{(1)^{\prime}}-\frac{1}{4}\partial^{i}w^{(1)}\partial_{k}D^{k}_{j}h^{(1)^{\prime}}\right]$
$-\frac{1}{a^{2}\kappa^{2}}\left(4\psi^{(1)}\delta^{ik}-2D^{ik}h^{(1)}\right)\left[\left(\partial_{j}\partial_{k}-\frac{1}{3}\nabla^{2}\delta_{jk}\right)\left(\phi^{(1)}-\psi^{(1)}\right.\right.$
$\left.\left.+2\mathcal{H}w^{(1)}+w^{(1)^{\prime}}\right)\right].$ (20)
Re-arranging and using the relation for the equation of state, $w=\bar{P}/\bar{\rho}$, and the background Section 3.1, we can find the wave-equation for the second-order tensor mode,
$h^{i(2)^{\prime\prime}}_{j}+2\mathcal{H}h^{i(2)^{\prime}}_{j}-\nabla^{2}h^{i(2)}_{j}=-4P^{li}_{jm}S^{m}_{l},$ (21)
where the source term is given in Appendix E. We decided to find this generic form of the source term Appendix E as this can be simplified to further understand the behavior in different gauges by imposing the gauge condition.
It is convenient to solve the Eq. 21 in Fourier space in which the source term will become a convolution of the first-order scalar perturbation at different wave-numbers. The second-order tensor mode can be written as
$h^{(2)}_{ij}(\mathbf{x},\tau)=\sum_{\lambda=+,\times}\dfrac{1}{(2\pi)^{3/2}}\int d^{3}\mathbf{k}\;e^{i\mathbf{k}\cdot\mathbf{x}}h^{(2)}_{\lambda}(\mathbf{k},\tau)e^{\lambda}_{ij}(\mathbf{k})\;,$ (22)
where $\lambda=+,\times$ denotes the two GW polarizations and $e^{\lambda}_{ij}(\mathbf{k})$ are the polarization tensors. The E.o.M for scalar-induced tensor modes can be written in the frequency domain as
$h^{(2)^{\prime\prime}}(\mathbf{k},\tau)+2\mathcal{H}h^{(2)^{\prime}}(\mathbf{k},\tau)+k^{2}h^{(2)}(\mathbf{k},\tau)=S_{\lambda}(\mathbf{k},\tau),$ (23)
where $S_{\lambda}(\tilde{\tau},\mathbf{k})$ includes the Fourier transformation of the source Appendix E and depends on the evolution of the scalar modes. Any scalar mode can be split into a transfer function $T(k\tau)$, which includes the time evolution of the modes after they re-enter the horizon, and the initial condition $\zeta_{\mathbf{k}}$, determined by inflation. For example, the scalar potential $\phi^{(1)}(\mathbf{k},\tau)$ can be split as
$\phi^{(1)}(\mathbf{k},\tau)=\frac{3(1+w)}{5+3w}T_{\phi}(k\tau)\zeta_{\mathbf{k}}\overset{w\rightarrow 1/3}{=}\frac{2}{3}T_{\phi}(k\tau)\zeta_{\mathbf{k}}.$ (24)
The primordial curvature fluctuations $\zeta_{\mathbf{k}}$ are characterised by the following power spectrum $\mathcal{P}_{\zeta}(k)$ as
$\langle\zeta_{\mathbf{k}}\zeta_{\mathbf{q}}\rangle=\delta^{(3)}(\mathbf{k}+\mathbf{q})\mathcal{P}_{\zeta}(k)=\dfrac{2\pi^{2}}{k^{3}}\delta^{(3)}(\mathbf{k}+\mathbf{q})\Delta^{2}_{\zeta}(k)\ ,$ (25)
where $\Delta^{2}_{\zeta}(k)$ is the dimensionless power spectrum. The source term can be written as
$S_{\lambda}(\mathbf{k},\tau)$ $=-4\mathbf{e}^{l}_{m}(\mathbf{k})S_{l}^{m}(\mathbf{k},\tau)$
$=4\int\frac{d^{3}\mathbf{q}}{(2\pi)^{3/2}}Q_{\lambda}(\mathbf{k},\mathbf{q})f(|\mathbf{k}-\mathbf{q}|,q,\tau)\zeta_{\mathbf{q}}\zeta_{\mathbf{k}-\mathbf{q}},$ (26)
where we define the polarization tensors, $Q_{\lambda}(\mathbf{k},\mathbf{q})\equiv\mathbf{e}^{l}_{\lambda m}(\mathbf{k})\mathbf{q}_{l}\mathbf{q}^{m}$ and $f(|\mathbf{k}-\mathbf{q}|,q,\tau)$ is the source function, which is found using the transfer functions of the scalar modes. In this review, we consider the production of SIGW to be when the scalar modes re-enter the causal horizon during the radiation-dominated (RD) era. Therefore, we will be performing the remaining calculations in the RD epoch, where the scale-factor behaves as $a\propto\tau$, resulting in $\mathcal{H}\propto\tau^{-1}$ and $w=1/3$. During the radiation epoch the dimensionless power spectrum for the tensor modes is defined by
$\langle h_{\lambda,\ \mathbf{k}}h_{\lambda^{\prime},\ \mathbf{q}}\rangle=\delta_{\lambda\lambda^{\prime}}\delta^{(3)}(\mathbf{k}+\mathbf{q})\dfrac{2\pi^{2}}{k^{3}}\Delta^{2}_{h}(k,\tau)\ ,$ (27)
where $\Delta^{2}_{h}(k,\tau)$ is
[…]
Notice the change to spherical momentum coordinates where we use dimensionless variables $u=\frac{|\mathbf{k}-\mathbf{q}|}{k}$ and $v=\frac{\mathbf{q}}{k}$. The function, $I(v,u,x)$ is known as the kernel and is found in relation to the source-term [10, 31, 32],
[…]
where the Green’s function $G_{\mathbf{k}}$ in radiation era is
$G_{\mathbf{k}}(\tau,\tilde{\tau})=\frac{1}{k}\sin(k\tau-k\tilde{\tau})\ .$ (30)
The fraction of GW energy density per logarithmic wavelength is defined by
$\Omega_{\rm GW}(k,\tau)\equiv\frac{\rho_{\rm GW}(k,\tau)}{\rho_{c}(\tau)}=\frac{1}{48}\left(\frac{k}{a(\tau)H(\tau)}\right)^{2}\sum_{\lambda=+,\times}\overline{{\Delta^{2}_{h,\lambda}(k,\tau)}}\ ,$ (31)
where $\rho_{c}(\tau)$ is the critical density and he averaged energy density of GWs is expressed as
$\rho_{GW}(\tau)=\frac{1}{16a^{2}\kappa^{2}(\tau)}\overline{\langle(\nabla h_{ij})^{2}\rangle}.$ (32)
The overline above the power spectrum is the oscillation average, derived from the average of the kernel.
4 Gauge-dependence of first-order metric perturbations
We now review how quantities in the metric change under coordinate transformation, also known as gauge-transformation. There are two approaches in studying gauge-transformation, the active and passive approach. We follow the approach taken by [33, 34, 18], which calculates the change in perturbations under coordinate change at a fixed physical point in spacetime. Later we look at the construction of the gauge-invariant variables.
4.1 First-order coordinate transformation of cosmological metric perturbations
The change in coordinate system for the time slicing and spatial threading is
$\tilde{\tau}$ $=\tau+\xi^{0},$ (33)
$\tilde{x}^{i}$ $=x^{i}+\partial_{i}\xi,$ (34)
where fixing $\xi^{0}$ determines the time-slicing and the spatial coordinate by $\xi$. With these transformations the perturbed line element, Section 2, can be written in terms of the transformed coordinate system. The scalar modes transform as
$\tilde{\phi}^{(1)}$ $=\phi^{(1)}-\mathcal{H}\xi^{0}-(\xi^{0})^{{}^{\prime}},$ (35)
$\tilde{\psi}^{(1)}$ $=\psi^{(1)}+\mathcal{H}\xi^{0},$ (36)
$\tilde{w}^{(1)}$ $=w^{(1)}+\xi^{0}-\xi^{{}^{\prime}},$ (37)
$\tilde{h}^{(1)}$ $=h^{(1)}-\xi.$ (38)
As mentioned in Section 2, first-order vector modes and tensor modes have negligible contribution to the second-order tensor modes and therefore not considered.
4.2 Gauge-invariant combination of first-order scalar perturbations
The gauge-invariant approach implies that quantities that are gauge dependent, in this case the scalar perturbations, can be constructed into gauge-invariant variables by defining a combination of perturbations variables which coincide with a particular gauge. This is obtained by fixing the coordinate transformations Eqs. 35, 36, 37 and 38. This is different from the gauge-independent quantities such as the first-order tensor perturbation, which is the same in all gauges. For that reason, one can find different gauge-invariant variables that correspond to different gauge choices and using these variables, the governing equations can be reformulated. Various gauges are chosen depending on the case which is being studied [35]. In the following section, we review gauges that have been used in literature in studying SIGWs.
4.2.1 Synchronous gauge
The synchronous gauge is widely used in numerical computations like those performed in Boltzmann codes (e.g. CLASS [36]). This gauge also known as time-orthogonal gauge and TT gauge, which is the typical choice of gauge for gravitational wave measurement and used in numerical studies. It is defined by
$g_{00}^{(1)}=g_{0i}^{(1)}=0,$ (39)
which allows us to set $\phi^{(1)}=\hat{w}_{i}^{(1)}=0$. This provides a simplification in the dynamical equations, as the cosmic time coordinates of FRW background coincide with the proper time of a observer at a fixed spatial coordinate, i.e. $d\tau=ad\tau$. One disadvantage with this gauge is that it is not possible to formulate a clear gauge-invariant variable to due residual gauge freedom. The gauge-transformation Eqs. 35 and 37 give the following
[…]
which shows that the time-slicing is not determined unambiguously and there remains two arbitrary scalar functions of the spatial coordinate $\mathcal{C}_{1}(x)$ and $\mathcal{\hat{C}}_{1}(x)$. Hence, there is no true gauge-invariant variable.
4.2.2 Poisson gauge
This gauge is defined by three conditions removing certain degrees of freedom; (i) vanishing shear, (ii) transverse condition on vector modes and (iii) transverse and traceless condition on tensor modes. Hence
$\nabla\cdot w=\nabla\cdot h=0.$ (42)
This gauge is also known as zero-shear gauge. Using this gauge, Bardeen [37] first proposed quantities that invariant under transformation. By studying the transformation equations, the corresponding gauge-invariant variables are known as the Bardeen potentials
$\Phi$ $=\phi^{(1)}-\mathcal{H}\sigma-\sigma^{\prime},$ (43)
$\Psi$ $=\psi^{(1)}+\mathcal{H}\sigma,$ (44)
where $\sigma^{(1)}\equiv h^{(1)^{\prime}}-w^{(1)}$ is the shear potential. In the framework of general perturbation theory, first-order vector and tensor perturbations remain present. However, by imposing the constraints $w_{i}=h_{ij}=0$, one recovers the Conformal Newtonian gauge. The Conformal Newtonian gauge is typically employed when vector and tensor perturbations are negligible, whereas the Poisson gauge is preferred in scenarios where vector perturbations play a significant role. In our analysis, since we focus on second-order tensor perturbations, we have already disregarded first-order vector and tensor contributions, leading to the same conditions as those in the Conformal Newtonian gauge.
The Poisson gauge is preferred due to its convenience in calculations on small scales. In the absence of anisotropic stress the scalar potentials are equal, i.e. $\Phi=\Psi$, and the governing field equations take a form close to the Newtonian ones.
4.2.3 Uniform curvature gauge
The uniform curvature gauge, also known as the spatially flat gauge, requires the spatial curvature to be unperturbed and homogeneous. This is accomplished by setting
$\psi^{(1)}=h^{(1)}=0,$ (45)
and the corresponding gauge-invariant variables are
$\tilde{\phi}^{(1)}$ $=\phi^{(1)}+\psi^{(1)}-\left(-\frac{\psi^{(1)}}{\mathcal{H}}\right)^{\prime},$ (46)
$\tilde{w}^{(1)}$ $=w^{(1)}-\frac{\psi^{(1)}}{\mathcal{H}}-h^{(1)^{\prime}}.$ (47)
This gauge is more convenient when studying observables during the inflationary period [35]. It is beneficial for analyzing scalar field dynamics and matter perturbations while eliminating the influence of metric fluctuations. It was also mentioned in [28], to being one of the “well-behaved” gauges on small scales for studying SIGWs.
5 Source function of SIGWs in different gauges
In this section, we analyze the scalar-scalar source term of SIGWs in the gauges mentioned above using two approaches: (i) deriving the source term in the synchronous and Poisson gauges via their gauge conditions, and (ii) employing gauge-invariant variables to study the source term in the uniform curvature gauge.
5.1 Scalar-scalar source in synchronous gauge
Using the definition for the shear potential, $\sigma^{(1)}$, and the constraint $\phi^{(1)}=w^{(1)}=0$, Appendix E simplifies to,
$S^{i}_{j,S}=$ $-\psi^{(1)^{\prime}}\partial^{i}\partial_{j}\sigma^{(1)}+2\partial^{i}\psi^{(1)^{\prime}}\partial_{j}\sigma^{(1)}+2\partial_{j}\psi^{(1)^{\prime}}\partial^{i}\sigma^{(1)}+\psi^{(1)}\partial^{i}\partial_{j}\psi^{(1)}$
$-\frac{1}{\mathcal{H}^{2}}\partial^{i}\psi^{(1)^{\prime}}\partial_{j}\psi^{(1)^{\prime}}+\partial^{i}\partial_{j}\sigma^{(1)}\nabla^{2}\sigma^{(1)}-\partial^{i}\partial_{k}\sigma^{(1)}\partial^{k}\partial_{j}\sigma^{(1)}.$ (48)
In Fourier space, the source function becomes
$f_{S}(|\mathbf{k}-\mathbf{q}|,q,\tau)=$ $\frac{4}{9}\left[T_{\psi}(q\tau)T_{\psi}(|\mathbf{k}-\mathbf{q}|\tau)-2T^{\prime}_{\psi}(q\tau)T_{\sigma}(|\mathbf{k}-\mathbf{q}|\tau)\right.$
$\left.-T_{\sigma}(q\tau)T^{\prime}_{\psi}(|\mathbf{k}-\mathbf{q}|\tau)+\tau^{2}T^{\prime}_{\psi}(q\tau)T^{\prime}_{\psi}(|\mathbf{k}-\mathbf{q}|\tau)\right.$
$\left.-(\mathbf{k}-\mathbf{q})T_{\sigma}(q\tau)T_{\sigma}(|\mathbf{k}-\mathbf{q}|\tau)-(\mathbf{k}\cdot\mathbf{q}-q^{2})T_{\sigma}(q\tau)T_{\sigma}(|\mathbf{k}-\mathbf{q}|\tau)\right].$ (49)
5.2 Scalar-scalar source in Poisson gauge
By using the gauge-condition $w^{(1)}=h^{(1)}=0$ and in the absence of anisotropic stress where $\phi^{(1)}=\psi^{(1)}$, we can simplify Appendix E in the Poisson gauge
$S^{i}_{j,P}$ $=4\phi^{(1)}\partial^{i}\partial_{j}\phi^{(1)}+\partial^{i}\phi^{(1)}\partial_{j}\phi^{(1)}$
$-\tau^{2}\left(\partial^{i}\phi^{(1)^{\prime}}\partial_{j}\phi^{(1)^{\prime}}+\tau^{-1}\partial^{i}\phi^{(1)^{\prime}}\partial_{j}\phi^{(1)}+\tau^{-1}\partial^{i}\phi^{(1)}\partial_{j}\phi^{(1)^{\prime}}\right).$ (50)
Therefore the source function in Fourier space is
$f_{P}(|\mathbf{k}-\mathbf{q}|,q,\tau)=$ $\frac{4}{9}\left[3T_{\phi}(q\tau)T_{\phi}(|\mathbf{k}-\mathbf{q}|\tau)+\tau^{2}T^{\prime}_{\phi}(q\tau)T^{\prime}_{\phi}(|\mathbf{k}-\mathbf{q}|\tau)\right.$
$\left.+\tau\left\{T_{\phi}(q\tau)T^{\prime}_{\phi}(|\mathbf{k}-\mathbf{q}|\tau)+T_{\phi}(q\tau)T^{\prime}_{\phi}(|\mathbf{k}-\mathbf{q}|\tau)\right\}\right].$ (51)
5.3 Scalar-scalar source in uniform curvature gauge
Next we will study the uniform curvature gauge where $\psi^{(1)}=h^{(1)}=0$. We can reduce Appendix E to
$S^{i}_{j,U}$ $=2\phi^{(1)}\partial^{i}\partial_{j}\phi^{(1)}+4\tau^{-1}\phi^{(1)}\partial^{i}\partial_{j}w^{(1)}+\phi^{(1)^{\prime}}\partial^{i}\partial_{j}w^{(1)}+2\phi^{(1)}\partial^{i}\partial_{j}w^{(1)^{\prime}}$
$+\nabla^{2}w^{(1)}\partial^{i}\partial_{j}w^{(1)}-\partial^{i}\partial_{k}w^{(1)}\partial^{k}\partial_{j}w^{(1)}.$ (52)
The standard way would be to find the solution for $\phi^{(1)}$ and $w^{(1)}$ and then later solve the E.o.M for SIGW. On the other hand, we have the gauge-invariant variables in the uniform curvature gauge, Eq. 46 and Eq. 47. We can therefore use this to our advantage. First we can make the same assumption of $\psi^{(1)}=\phi^{(1)}$ as done in the Poisson gauge and re-write the gauge-invariant variables as
$\tilde{\phi}^{(1)}$ $=2\phi^{(1)}-\left(-\frac{\phi^{(1)}}{\mathcal{H}}\right)^{\prime},$ (53)
$\tilde{w}^{(1)}$ $=-\frac{\phi^{(1)}}{\mathcal{H}}.$ (54)
As done in the Poisson gauge, the source term can be written in Fourier space in terms of the gauge-invariant variables and the transfer functions for the scalar potential, $\phi^{(1)}$,
$f_{U}(|\mathbf{k}-\mathbf{q}|,q,\tau)=$ $\frac{4}{9}\left[4\tau T_{\phi}(q\tau)T^{\prime}_{\phi}(|\mathbf{k}-\mathbf{q}|\tau)-\tau^{2}T^{\prime\prime}_{\phi}(q\tau)T_{\phi}(|\mathbf{k}-\mathbf{q}|\tau)\right.$
$\left.-8\tau T^{\prime}_{\phi}(q\tau)T_{\phi}(|\mathbf{k}-\mathbf{q}|\tau)-\tau^{2}(\mathbf{k}-\mathbf{q})T_{\phi}(q\tau)T_{\phi}(|\mathbf{k}-\mathbf{q}|\tau)\right.$
$\left.-\tau^{2}(\mathbf{k}\cdot\mathbf{q}-q^{2})T_{\phi}(q\tau)T_{\phi}(|\mathbf{k}-\mathbf{q}|\tau)\right].$ (55)
This means that we will use the solution for the scalar-potential in the Poisson gauge to evaluate the source function in the uniform curvature gauge.
6 Kernel in various gauges
In this section we derive the analytical expressions of the source function and later perform numerical integration of the kernel in the various gauges. As seen in Section 5, the source function in the synchronous gauge, Section 5.1, is found in terms of the transfer functions for the scalar modes $T_{\psi}$ and $T_{\sigma}$. We study the behavior in synchronous gauge, following the method outlined in [27], by deriving the linear solutions for the scalar modes and subsequently simplify the source term. On the other hand, the source function in the Poisson gauge, Section 5.2, and uniform gauge, Section 5.3, are written in terms of the transfer function for the scalar potential $\phi^{(1)}$. We can repeat the same analysis of the first-order scalar potential in the Poisson gauge. Using the transfer function, $T_{\phi}$, we are able to compute both in the Poisson and uniform gauge.
6.1 Linear solutions in synchronous gauge
We apply the definition $\phi^{(1)}=w^{(1)}=0$ to the first-order equations Section 3.2Eq. 17 and using the definition for shear, $\sigma^{(1)}$, the equations of motion for $\psi^{(1)}$ and $\sigma^{(1)}$ is obtained
$2\mathcal{H}\sigma^{(1)}+\sigma^{(1)^{\prime}}+\psi^{(1)}$ $=0,$ (56)
$6\psi^{(1)^{\prime\prime}}+2\mathcal{H}\left(9\psi^{(1)^{\prime}}-4\nabla^{2}\sigma^{(1)}\right)-3\nabla^{2}\sigma^{(1)^{\prime}}-5\nabla^{2}\psi^{(1)}$ $=0.$ (57)
The solutions are given as
$T_{\psi}(k,\tau)=9\left(\frac{\sin\left(\frac{k\tau}{\sqrt{3}}\right)-k\tau}{(k\tau)^{2}}\right),$ (58)
$T_{\sigma}(k,\tau)=9\left(\frac{1-\cos\left(\frac{k\tau}{\sqrt{3}}\right)}{(k\tau)^{2}}\right).\$ (59)
6.2 Linear solutions in Poisson gauge
As mentioned in the previous sections, this gauge is defined by $w^{(1)}=h^{(1)}=0$. In the first place we see that the linear-order trace-free part of the spatial Einstein field equation, Eq. 17 provides us with the following constraint equation
$\left(\partial^{i}\partial_{j}-\frac{1}{3}\nabla^{2}\delta^{i}_{j}\right)\left(\phi^{(1)}-\psi^{(1)}\right)=\kappa^{2}\pi^{(1)i}_{j}\ ,$ (60)
The source anisotropic stress comes from free-streaming neutrinos which is minimal and therefore we can neglect $\pi^{(1)i}_{j}$ and set $\psi^{(1)}=\phi^{(1)}$. The linear-order solution for the scalar potential is found using the first-order evolution equations, Section 3.2 and Section 3.2
$\phi^{(1)^{\prime\prime}}+3\mathcal{H}\left(1+c_{s}^{2}\right)\phi^{(1)^{\prime}}+\left(2\mathcal{H}^{\prime}+\left(1+3c_{s}^{2}\right)\mathcal{H}^{2}+c_{s}^{2}k^{2}\right)\phi^{(1)}=0\ ,$ (61)
The general solution in RD era is found as Bessel function,
$T_{\phi}(k,\tau)=\frac{9}{(k\tau)^{2}}\left[\frac{\sqrt{3}}{k\tau}\sin\left(\frac{k\tau}{\sqrt{3}}\right)-\cos\left(\frac{k\tau}{\sqrt{3}}\right)\right].$ (62)
6.3 Analytical expression of the source function
Here we present the analytical expression of the kernel in all three gauges. Starting with the synchronous gauge. We substitute the transfer functions Eq. 58Eq. 59 into Section 5.1 to find
$f_{S}(u,v,\tau)=\frac{2}{u^{3}v^{3}x^{5}}$ $\left(-9uvx\left(x^{2}\left(u^{2}-v^{2}+1\right)+8u+4v-10\right)\right.$
$\left.+3\sin\left(\frac{vx}{\sqrt{3}}\right)\left(\sqrt{3}u\left(x^{2}\left(3\left(u^{2}+1\right)+(4u-7)v^{2}\right)+12v\right)\right.\right.$
$\left.\left.+x\left(u^{2}\left(2v\left(vx^{2}-3\right)-9\right)-12uv^{2}+9v^{2}-9\right)\sin\left(\frac{ux}{\sqrt{3}}\right)\right)\right.$
$\left.+3\sqrt{3}v\left(x^{2}\left(u^{2}(2v-1)-3v^{2}+3\right)+24u\right)\sin\left(\frac{ux}{\sqrt{3}}\right)\right.$
$\left.+6uv\left(2\sqrt{3}\left(ux^{2}-6\right)\sin\left(\frac{ux}{\sqrt{3}}\right)+3(4u-5)x\right)\cos\left(\frac{vx}{\sqrt{3}}\right)\right.$
$\left.+6uv\cos\left(\frac{ux}{\sqrt{3}}\right)\left(2\sqrt{3}\left(vx^{2}-3\right)\sin\left(\frac{vx}{\sqrt{3}}\right)\right.\right.$
$\left.\left.+3(2v-5)x+15x\cos\left(\frac{vx}{\sqrt{3}}\right)\right)\right).$ (63)
The source function for Poisson and uniform curvature gauge are found by substituting Eq. 62 into Section 5.2 and Section 5.3 respectively,
$f_{P}(u,v,\tau)=\frac{12}{u^{3}v^{3}x^{6}}$ $\left(\sin\left(\frac{ux}{\sqrt{3}}\right)\left(\left(u^{2}x^{2}\left(v^{2}x^{2}-6\right)-6v^{2}x^{2}+54\right)\sin\left(\frac{vx}{\sqrt{3}}\right)\right.\right.$
$\left.\left.+2\sqrt{3}vx\left(u^{2}x^{2}-9\right)\cos\left(\frac{vx}{\sqrt{3}}\right)\right)\right.$
$\left.+2ux\cos\left(\frac{ux}{\sqrt{3}}\right)\left(\sqrt{3}\left(v^{2}x^{2}-9\right)\sin\left(\frac{vx}{\sqrt{3}}\right)\right.\right.$
$\left.\left.+9vx\cos\left(\frac{vx}{\sqrt{3}}\right)\right)\right),$ (64)
$f_{U}(u,v,\tau)=-\frac{6}{u^{3}v^{3}x^{4}}$ $\left(\sin\left(\frac{ux}{\sqrt{3}}\right)\left(3\left(-5u^{2}+3v^{2}+3\right)\sin\left(\frac{vx}{\sqrt{3}}\right)\right.\right.$
$\left.\left.+\sqrt{3}vx\left(5u^{2}+5v^{2}-3\right)\cos\left(\frac{vx}{\sqrt{3}}\right)\right)\right.$
$\left.+ux\cos\left(\frac{ux}{\sqrt{3}}\right)\left(vx\left(3u^{2}-5v^{2}+3\right)\cos\left(\frac{vx}{\sqrt{3}}\right)\right.\right.$
$\left.\left.-3\sqrt{3}\left(u^{2}+v^{2}+1\right)\sin\left(\frac{vx}{\sqrt{3}}\right)\right)\right).$ (65)
6.4 Numerical integration of the kernel
We numerically evaluate the kernel, Eq. 29, in the three different gauges using the source functions Section 6.3, Section 6.3 and Section 6.3. By fixing a set of values for $u$ and $v$, the numerical integration is presented in Fig. 1. We show the comparison of the numerically integrated kernel in all three gauges with the oscillation average of the kernel function, Eq.(26) from [32] which was found in the conformal Newtonian gauge. In addition, we present residual between our numerically integrated kernel and the analytical solution [32].
Figure 1: The kernel function $I^{2}(v=1,u=1,x)$ as a function of $x=k\tau$. In the left panel we show the oscillation average of the kernel function (red line) taken from Eq.(26) of [32] compared to numerical integration of the kernel in the synchronous (green line), Poisson (blue line) and uniform curvature (orange line) gauges. In the right panel, we plot the residual between the numerically integrated kernel in three gauges and the analytical integration.
We observe that the numerical results from the different gauge choices closely follow the analytical solution. The kernel values for the Poisson, uniform curvature, and synchronous gauges exhibit minimal variation, indicating that any residual gauge effects are small. The oscillatory structure seen in the analytic oscillation average seems to align with the numerically integrated kernels. From the panel in the right, we notice the residuals appear small across all gauges, confirming that the gauge choices do not significantly affect the final gravitational wave kernel. The synchronous gauge may introduce this numerical discrepancies due to its residual gauge freedom. Even if there is a minimal discrepancy, the variation in behavior in all three gauges decreases as $x\gg 1$, when entering deep in the horizon. This suggests that despite gauge differences, the kernel remains consistent, confirming gauge invariance at the level of physical observables.
7 Conclusion
In this review, we have revisited the issue of gauge-dependence of “scalar-induced” gravitational waves, which are second-order gravitational waves produced by first-order scalars. We approach this issue by addressing the gauge-invariant approach which has already been discussed in previous studies [20, 22, 38, 27, 28]. We start by studying the metric perturbation up to second-order without imposing a gauge and retrieve the scalar-scalar source term for the second-order tensor perturbations. Next we revise the gauge-invariant approach by studying the gauge-transformation properties and deriving the gauge-invariant variables for different gauges; Poisson, uniform curvature and synchronous. We adopt two approaches in finding the behavior in the different gauge. For the synchronous gauge, we find the subsequent first-order solutions for the perturbations in that gauge and compute the kernel. For the analysis in Poisson and uniform curvature gauge, we use the gauge invariant method by initially finding the first-order solution in the Poisson gauge and then following similar procedure to [38], we use the gauge-invariant variables found in Section 4.2.3 to then find the kernel in the uniform curvature gauge. Our computation has been done numerically in the radiation-dominant case, with $w=1/3$.
The results confirm that the kernel gravitational waves remains robust across different gauge choices similar to findings from [21, 39]. The numerical integration in the Poisson gauge successfully reproduces the expected analytical behavior from [32], validating the computational approach, since the Poisson gauge reduces to the Conformal Newtonian gauge in the absence of vector perturbations. As discussed previously, the Poisson gauge is often preferred for such calculations due to its direct relation to Newtonian-like potentials, while the synchronous gauge can introduce coordinate artifacts due to its residual gauge freedom.
[…]
Acknowledgments I would like to thank Sabino Matarrese and Andrea Maselli for their insightful comments and discussions. My gratitude also goes to Marisol Traforetti for her assistance with the computation and numerical integration of the kernels.
Appendix A Ricci tensor
The Ricci tensor, which is a contraction of the Riemann tensor, can be written in terms of the connection coefficient
$R_{\mu\nu}=\partial_{\alpha}\Gamma^{\alpha}_{\mu\nu}-\partial_{\nu}\Gamma^{\alpha}_{\nu\alpha}+\Gamma^{\alpha}_{\sigma\alpha}\Gamma^{\sigma}_{\mu\nu}-\Gamma^{\alpha}_{\sigma\nu}\Gamma^{\sigma}_{\mu\alpha},$ (66)
where the connection coefficients is defined as
$\Gamma^{\alpha}_{\beta\gamma}=\frac{1}{2}g^{\alpha\sigma}\left(\frac{\partial g_{\sigma\gamma}}{\partial x^{\beta}}+\frac{\partial g_{\beta\sigma}}{\partial x^{\gamma}}-\frac{\partial g_{\beta\gamma}}{\partial x^{\sigma}}\right).$ (67)
The perturbed components are,
$R^{(0)}_{00}$ $=-3\frac{a^{\prime\prime}}{a}+3\left(\frac{a^{\prime}}{a}\right)^{2},$ (68)
$R^{(0)}_{0i}$ $=0,$ (69)
$R^{(0)}_{ij}$ $=\left[\frac{a^{\prime\prime}}{a}+\left(\frac{a^{\prime}}{a}\right)^{2}\right]\delta^{i}_{j}.$ (70)
$R^{(1)}_{00}$ $=\nabla^{2}\phi^{(1)}+3\frac{a^{\prime}}{a}\phi^{(1)^{\prime}}+3\psi^{(1)^{\prime\prime}}+3\frac{a^{\prime}}{a}\psi^{(1)^{\prime}}+\frac{a^{\prime}}{a}\nabla^{2}w^{(1)}+\nabla^{2}w^{(1)^{\prime}},$ (71)
$R^{(1)}_{0i}$ $=2\frac{a^{\prime}}{a}\partial_{i}\phi^{(1)}+2\partial_{i}\psi^{(1)^{\prime}}+\left(\frac{a^{\prime\prime}}{a}+\left(\frac{a^{\prime}}{a}\right)^{2}\right)\partial_{i}w^{(1)}+\frac{1}{2}\partial_{k}D^{k}_{i}h^{(1)},$ (72)
$R^{(1)}_{ij}$ $=\left[-2\left(\frac{a^{\prime\prime}}{a}+\frac{a^{\prime}}{a}\right)^{2}\phi^{(1)}-\frac{a^{\prime}}{a}\phi^{(1)^{\prime}}-2\left(\frac{a^{\prime\prime}}{a}+\frac{a^{\prime}}{a}\right)^{2}\psi^{(1)}-5\frac{a^{\prime}}{a}\psi^{(1)^{\prime}}\right.$
$\left.-\psi^{(1)^{\prime\prime}}+\partial^{k}\partial_{k}\psi^{(1)}-\frac{a^{\prime}}{a}\nabla^{2}w^{(1)}\right]\delta^{i}_{j}+\partial_{i}\partial_{j}\psi^{(1)}-\partial_{i}\partial_{j}\phi^{(1)}$
$-2\frac{a^{\prime}}{a}\partial_{i}\partial_{j}w^{(1)}-\partial_{i}\partial_{j}w^{(1)^{\prime}}+\left(\frac{a^{\prime\prime}}{a}+\frac{a^{\prime}}{a}\right)^{2}D_{ij}h^{(1)}+\frac{a^{\prime}}{a}D_{ij}h^{(1)^{\prime}}$
$+\frac{1}{2}D_{ij}h^{(1)^{\prime\prime}}+\frac{1}{2}\partial_{i}\partial_{k}D^{k}_{j}h^{(1)}+\frac{1}{2}\partial_{j}\partial_{j}D^{k}_{i}h^{(1)}-\frac{1}{2}\nabla^{2}D_{ij}h^{(1)}.$ (73)
$R^{(2)}_{00}$ $=\frac{3}{2}\frac{a^{\prime}}{a}\phi^{(2)^{\prime}}+\frac{1}{2}\nabla^{2}\phi^{(2)}+\frac{3}{2}\frac{a^{\prime}}{a}\psi^{(2)^{\prime}}+\frac{3}{2}\psi^{(2)^{\prime\prime}}+\frac{1}{2}\frac{a^{\prime}}{a}\nabla^{2}w^{(2)}+\frac{1}{2}\nabla^{2}w^{(2)^{\prime}}$
$-6\frac{a^{\prime}}{a}\phi^{(1)}\phi^{(1)^{\prime}}-\partial^{k}\phi^{(1)}\partial_{k}\phi^{(1)}-3\phi^{(1)^{\prime}}\psi^{(1)^{\prime}}+2\psi^{(1)}\nabla^{2}\phi^{(1)}-\partial_{k}\psi^{(1)}\partial^{k}\phi^{(1)}$
$+6\frac{a^{\prime}}{a}\psi^{(1)}\psi^{(1)^{\prime}}+6\psi^{(1)}\psi^{(1)^{\prime\prime}}+3(\psi^{(1)^{\prime}})^{2}-\phi^{(1)^{\prime}}\nabla^{2}w^{(1)}+\frac{a^{\prime}}{a}\partial_{k}\phi^{(1)}\partial^{k}w^{(1)}$
$-\frac{a^{\prime}}{a}\partial_{k}\psi^{(1)}\partial^{k}w^{(1)}+2\frac{a^{\prime}}{a}\psi^{(1)}\nabla^{2}w^{(1)}-\partial_{k}\psi^{(1)}\partial^{k}w^{(1)^{\prime}}+2\psi^{(1)}\nabla^{2}w^{(1)^{\prime}}$
$+\left(\frac{a^{\prime\prime}}{a}+\left(\frac{a^{\prime}}{a}\right)^{2}\right)\partial_{k}w^{(1)}\partial^{k}w^{(1)}+3\frac{a^{\prime}}{a}\partial_{k}w^{(1)^{\prime}}\partial^{k}w^{(1)}-\partial_{k}\phi^{(1)}\partial_{m}D^{km}h^{(1)}$
$-\partial_{k}\partial_{m}\phi^{(1)}D^{km}h^{(1)}-\frac{a^{\prime}}{a}\partial_{k}w^{(1)}\partial_{m}D^{km}h^{(1)}-\frac{a^{\prime}}{a}\partial_{k}\partial_{m}w^{(1)}D^{km}h^{(1)}$
$-\partial_{k}w^{(1)^{\prime}}\partial_{m}D^{km}h^{(1)}-\partial_{k}\partial_{m}w^{(1)^{\prime}}D^{km}h^{(1)}-\frac{1}{2}\frac{a^{\prime}}{a}D^{km}h^{(1)}D_{km}h^{(1)^{\prime}}$
$+\frac{1}{4}D^{km}h^{(1)^{\prime}}D_{km}h^{(1)^{\prime}}+\frac{1}{2}D^{km}h^{(1)}D_{km}h^{(1)^{\prime\prime}},$ (74)
$R^{(2)}_{0i}$ $=\frac{a^{\prime}}{a}\partial_{i}\phi^{(2)}+\partial_{i}\psi^{(2)^{\prime}}-\frac{1}{4}\nabla^{2}w_{i}^{(2)}+\frac{1}{2}\left[\frac{a^{\prime\prime}}{a}+\left(\frac{a^{\prime}}{a}\right)^{2}\right]w_{i}^{(2)}$
$+\frac{1}{2}\left[\frac{a^{\prime\prime}}{a}+\left(\frac{a^{\prime}}{a}\right)^{2}\right]\partial_{i}w^{(2)}-\frac{1}{2}\nabla^{2}h^{(2)^{\prime}}_{i}+\frac{1}{4}\partial_{k}D^{k}_{i}h^{(2)^{\prime}}$
$-4\frac{a^{\prime}}{a}\phi^{(1)}\partial_{i}\phi^{(1)}-2\psi^{(1)^{\prime}}\partial_{i}\phi^{(1)}+4\psi^{(1)^{\prime}}\partial_{i}\psi^{(1)}+4\psi^{(1)}\partial_{i}\psi^{(1)^{\prime}}$
$-2\left[\frac{a^{\prime\prime}}{a}+\left(\frac{a^{\prime}}{a}\right)^{2}\right]\phi^{(1)}\partial_{i}w^{(1)}-\partial_{i}\phi^{(1)}\nabla^{2}w^{(1)}$
$-\partial_{i}\partial_{k}\phi^{(1)}\partial^{k}w^{(1)}+\partial^{k}\phi^{(1)}\partial_{i}\partial_{k}w^{(1)}-\frac{a^{\prime}}{a}\phi^{(1)^{\prime}}\partial_{i}w^{(1)}$
$-5\frac{a^{\prime}}{a}\psi^{(1)^{\prime}}\partial_{i}w^{(1)}-\psi^{(1)^{\prime\prime}}\partial_{i}w^{(1)}-\frac{a^{\prime}}{a}\partial_{i}w^{(1)}\nabla^{2}w^{(1)}-\partial^{k}w^{(1)}\partial_{i}\partial_{k}w^{(1)^{\prime}}$
$-\frac{1}{2}\partial_{k}\phi^{(1)}D^{k}_{i}h^{(1)^{\prime}}+\frac{1}{2}\partial_{k}\psi^{(1)}D^{k}_{i}h^{(1)^{\prime}}+\psi^{(1)}\partial_{k}D^{k}_{i}h^{(1)^{\prime}}$
$+\partial_{k}\psi^{(1)^{\prime}}D^{k}_{i}h^{(1)}+\psi^{(1)^{\prime}}\partial_{k}D^{k}_{i}h^{(1)}+\frac{a^{\prime}}{a}\partial^{k}w^{(1)}D_{ik}h^{(1)^{\prime}}$
$+\frac{1}{2}\partial^{k}w^{(1)}D_{ik}h^{(1)^{\prime\prime}}-\frac{1}{2}D^{km}h^{(1)}\partial_{k}D_{mi}h^{(1)^{\prime}}+\frac{1}{2}D^{km}h^{(1)}\partial_{i}D_{km}h^{(1)^{\prime}}$
$-\frac{1}{2}\partial_{k}D^{km}h^{(1)}D_{mi}h^{(1)^{\prime}}+\frac{1}{4}\partial_{i}D^{km}h^{(1)}D_{km}h^{(1)^{\prime}},$ (75)
$R^{(2)}_{ij}$ $=\left[-\left(\frac{a^{\prime\prime}}{a}+\left(\frac{a^{\prime}}{a}\right)^{2}\right)\phi^{(2)}-\frac{1}{2}\frac{a^{\prime}}{a}\phi^{(2)^{\prime}}-\left(\frac{a^{\prime\prime}}{a}+\left(\frac{a^{\prime}}{a}\right)^{2}\right)\psi^{(2)}-\frac{5}{2}\frac{a^{\prime}}{a}\psi^{(2)^{\prime}}\right.$
$\left.-\frac{1}{2}\psi^{(2)^{\prime\prime}}+\frac{1}{2}\nabla^{2}\psi^{(2)}-\frac{1}{2}\frac{a^{\prime}}{a}\nabla^{2}w^{(2)}+4\left(\frac{a^{\prime\prime}}{a}+\left(\frac{a^{\prime}}{a}\right)^{2}\right)(\phi^{(1)})^{2}\right.$
$\left.+4\frac{a^{\prime}}{a}\phi^{(1)}\phi^{(1)^{\prime}}+19\frac{a^{\prime}}{a}\phi^{(1)}\psi^{(1)^{\prime}}+2\frac{a^{\prime}}{a}\phi^{(1)^{\prime}}\psi^{(1)}+\phi^{(1)^{\prime}}\psi^{(1)^{\prime}}+2\phi^{(1)}\psi^{(1)^{\prime\prime}}\right.$
$\left.4\left(\frac{a^{\prime\prime}}{a}+\left(\frac{a^{\prime}}{a}\right)^{2}\right)\phi^{(1)}\psi^{(1)}+\partial_{k}\psi^{(1)}\partial^{k}\phi^{(1)}+(\psi^{(1)^{\prime}})^{2}+\partial_{k}\psi^{(1)}\partial^{k}\psi^{(1)}\right.$
$\left.+2\psi^{(1)}\nabla^{2}\psi^{(1)}+\frac{a^{\prime}}{a}\partial_{k}\phi^{(1)}\partial^{k}w^{(1)}+2\frac{a^{\prime}}{a}\phi^{(1)}\nabla^{2}w^{(1)}\right.$
$\left.+3\frac{a^{\prime}}{a}\partial^{k}\psi^{(1)}\partial_{k}w^{(1)}+\partial_{k}\psi^{(1)}\partial^{k}w^{(1)^{\prime}}+2\partial_{k}\psi^{(1)^{\prime}}\partial^{k}w^{(1)}\right.$
$\left.+\psi^{(1)^{\prime}}\nabla^{2}w^{(1)}-\partial_{m}\psi^{(1)}\partial_{k}D^{km}h^{(1)}-\partial_{k}\partial_{m}\psi^{(1)}D^{km}h^{(1)}\right.$
$\left.+\frac{a^{\prime}}{a}\partial_{m}\partial^{k}w^{(1)}D^{m}_{k}h^{(1)}+\frac{a^{\prime}}{a}\partial^{k}w^{(1)}\partial_{m}D^{m}_{k}h^{(1)}-\frac{1}{2}\frac{a^{\prime}}{a}D^{mk}h^{(1)}D_{km}h^{(1)^{\prime}}\right]\delta_{ij}$
$-\frac{1}{2}\partial_{i}\partial_{j}\phi^{(2)}+\frac{1}{2}\partial_{i}\partial_{j}\psi^{(2)}-\frac{a^{\prime}}{a}\partial_{i}\partial_{j}w^{(2)}-\frac{1}{2}\partial_{i}\partial_{j}w^{(2)^{\prime}}$
$-\frac{1}{2}\frac{a^{\prime}}{a}(\partial_{j}w_{i}^{(2)}+\partial_{i}w_{j}^{(2)})-\frac{1}{4}(\partial_{j}w_{i}^{(2)^{\prime}}+\partial_{i}w_{j}^{(2)^{\prime}})$
$+\frac{1}{2}\left[\frac{a^{\prime\prime}}{a}+\left(\frac{a^{\prime}}{a}\right)^{2}\right]\left(D_{ij}h^{(2)})+\partial_{i}h_{j}^{(2)}+\partial_{j}h^{(2)}_{i}+h_{ij}^{(2)}\right)$
$+\frac{1}{2}\frac{a^{\prime}}{a}\left(D_{ij}h^{(2)^{\prime}})+\partial_{i}h_{j}^{(2)^{\prime}}+\partial_{j}h^{(2)^{\prime}}_{i}+h_{ij}^{(2)^{\prime}}\right)$
$+\frac{1}{4}\left(D_{ij}h^{(2)^{\prime\prime}})+\partial_{i}h_{j}^{(2)^{\prime\prime}}+\partial_{j}h^{(2)^{\prime\prime}}_{i}+h_{ij}^{(2)^{\prime\prime}}\right)$
$+\frac{1}{2}\partial_{k}\partial_{i}D^{k}_{j}h^{(2)}-\frac{1}{4}\nabla^{2}h_{ij}^{(2)}-\frac{1}{4}\nabla^{2}D_{ij}h^{(2)}$
$+\partial_{i}\phi^{(1)}\partial_{j}\phi^{(1)}+2\phi^{(1)}\partial_{i}\partial_{j}\phi^{(1)}-\partial_{i}\phi^{(1)}\partial_{j}\psi^{(1)}-\partial_{i}\psi^{(1)}\partial_{j}\phi^{(1)}+3\partial_{i}\psi^{(1)}\partial_{j}\psi^{(1)}$
$+2\psi^{(1)}\partial_{i}\partial_{j}\psi^{(1)}+4\frac{a^{\prime}}{a}\phi^{(1)}\partial_{i}\partial_{j}w^{(1)}+\phi^{(1)^{\prime}}\partial_{i}\partial_{j}w^{(1)}+2\phi^{(1)}\partial_{i}\partial_{j}w^{(1)^{\prime}}$
$-2\frac{a^{\prime}}{a}\partial_{i}\psi^{(1)}\partial_{j}w^{(1)}-2\frac{a^{\prime}}{a}\partial_{j}\psi^{(1)}\partial_{i}w^{(1)}-\partial_{i}\psi^{(1)^{\prime}}\partial_{j}w^{(1)}-\partial_{j}\psi^{(1)^{\prime}}\partial_{i}w^{(1)}$
$-\partial_{j}\psi^{(1)}\partial_{i}w^{(1)^{\prime}}-\partial_{i}\psi^{(1)}\partial_{j}w^{(1)^{\prime}}+\psi^{(1)^{\prime}}\partial_{i}\partial_{j}w^{(1)}+\partial_{i}\partial_{j}w^{(1)}\nabla^{2}w^{(1)}$
$-\partial_{i}\partial_{k}w^{(1)}\partial_{j}\partial^{k}w^{(1)}-2\left[\left(\frac{a^{\prime}}{a}\right)^{2}+\frac{a^{\prime\prime}}{a}\right]\phi^{(1)}D_{ij}h^{(1)}-2\frac{a^{\prime}}{a}\phi^{(1)}D_{ij}h^{(1)^{\prime}}$
$-\frac{a^{\prime}}{a}\phi^{(1)^{\prime}}D_{ij}h^{(1)}-\frac{1}{2}\phi^{(1)^{\prime}}D_{ij}h^{(1)^{\prime}}\phi^{(1)}D_{ij}h^{(1)^{\prime\prime}}+\frac{1}{2}\partial_{k}\phi^{(1)}\partial_{i}D^{k}_{j}h^{(1)}$
$+\frac{1}{2}\partial_{k}\phi^{(1)}\partial_{j}D^{k}_{i}h^{(1)}-\frac{1}{2}\partial_{k}\phi^{(1)}\partial^{k}D_{ij}h^{(1)}+\frac{1}{2}\partial_{k}\psi^{(1)}\partial_{i}D^{k}_{j}h^{(1)}+\frac{1}{2}\partial_{k}\psi^{(1)}\partial_{j}D^{k}_{i}h^{(1)}$
$-\frac{3}{2}\partial_{k}\psi^{(1)}\partial^{k}D_{ij}h^{(1)}+\psi^{(1)}\partial_{k}\partial_{i}D^{k}_{j}h^{(1)}+\psi^{(1)}\partial_{k}\partial_{j}D^{k}_{i}h^{(1)}-\psi^{(1)}\nabla^{2}D_{ij}h^{(1)}$
$+\partial_{i}\psi^{(1)}\partial_{k}D^{k}_{j}h^{(1)}+\partial_{j}\psi^{(1)}\partial_{k}D^{k}_{i}h^{(1)}+\partial_{i}\partial_{k}\psi^{(1)}D^{k}_{j}h^{(1)}+\partial_{j}\partial_{k}\psi^{(1)}D^{k}_{i}h^{(1)}$
$-3\frac{a^{\prime}}{a}\psi^{(1)^{\prime}}D_{ij}h^{(1)}+\frac{1}{2}\psi^{(1)^{\prime}}D_{ij}h^{(1)^{\prime}}+\frac{a^{\prime}}{a}\partial^{k}w^{(1)}\partial_{i}D_{kj}h^{(1)}+\frac{a^{\prime}}{a}\partial^{k}w^{(1)}\partial_{j}D_{ki}h^{(1)}$
$-\frac{a^{\prime}}{a}\partial^{k}w^{(1)}\partial_{k}D_{ij}h^{(1)}-\frac{a^{\prime}}{a}\nabla^{2}w^{(1)}D_{ij}h^{(1)}+\frac{1}{2}\partial^{k}w^{(1)^{\prime}}\partial_{i}D_{kj}h^{(1)}$
$+\frac{1}{2}\partial^{k}w^{(1)^{\prime}}\partial_{j}D_{ki}h^{(1)}-\frac{1}{2}\partial^{k}w^{(1)^{\prime}}\partial_{k}D_{ij}h^{(1)}+\frac{1}{2}\partial^{k}w^{(1)}\partial_{i}D_{kj}h^{(1)^{\prime}}$
$+\frac{1}{2}\partial^{k}w^{(1)}\partial_{j}D_{ki}h^{(1)^{\prime}}-\partial^{k}w^{(1)}\partial_{k}D_{ij}h^{(1)^{\prime}}+\frac{1}{2}\partial_{k}\partial_{i}w^{(1)}D^{k}_{j}h^{(1)^{\prime}}$
$+\frac{1}{2}\partial_{k}\partial_{j}w^{(1)}D^{k}_{i}h^{(1)^{\prime}}-\frac{1}{2}\nabla^{2}w^{(1)}D_{ij}h^{(1)^{\prime}}-\frac{1}{2}\partial_{i}D_{mj}h^{(1)}\partial_{k}D^{km}h^{(1)}$
$-\frac{1}{2}\partial_{j}D_{mi}h^{(1)}\partial_{k}D^{km}h^{(1)}+\frac{1}{2}\partial_{m}D_{ij}h^{(1)}\partial_{k}D^{km}h^{(1)}-\frac{1}{2}\partial_{i}\partial_{k}D_{mj}h^{(1)}D^{km}h^{(1)}$
$-\frac{1}{2}\partial_{j}\partial_{k}D_{mi}h^{(1)}D^{km}h^{(1)}+\frac{1}{2}\partial_{m}\partial_{k}D_{ij}h^{(1)}D^{km}h^{(1)}+\frac{1}{2}\partial_{i}\partial_{j}D_{km}h^{(1)}D^{km}h^{(1)}$
$+\frac{1}{4}\partial_{j}D_{km}h^{(1)}\partial_{i}D^{km}h^{(1)}+\frac{1}{2}\partial_{m}D_{ik}h^{(1)}\partial^{m}D^{k}_{j}h^{(1)}$
$-\frac{1}{2}\partial_{m}D_{ik}h^{(1)}\partial^{k}D^{m}_{j}h^{(1)}-\frac{1}{2}D^{k}_{i}h^{(1)^{\prime}}D_{kj}h^{(1)^{\prime}}.$ (76)
Appendix B Ricci scalar
The Ricci scalar is a contraction of the Ricci tensor
$R=R^{\mu}_{\mu}$ (77)
The components are expanded to the following perturbations,
$R^{(0)}=\frac{6}{a^{2}}\frac{a^{\prime\prime}}{a^{\prime}},$ (78)
$R^{(1)}$ $=\frac{1}{a^{2}}\left(-2\nabla^{2}\phi^{(1)}-6\psi^{(1)^{\prime\prime}}-6\frac{a^{\prime}}{a}\phi^{(1)^{\prime}}-18\frac{a^{\prime}}{a}\psi^{(1)^{\prime}}-12\frac{a^{\prime\prime}}{a}\phi^{(1)}+4\nabla^{2}\psi^{(1)}\right.$
$\left.-6\frac{a^{\prime}}{a}\nabla^{2}w^{(1)}-2\nabla^{2}w^{(1)^{\prime}}+\partial^{k}\partial_{m}D^{m}_{k}h^{(1)}\right),$ (79)
$R^{(2)}$ $=\frac{1}{a^{2}}\left(-\nabla^{2}\phi^{(2)}-3\frac{a^{\prime}}{a}\phi^{(2)^{\prime}}-6\frac{a^{\prime\prime}}{a}\phi^{(2)}+2\nabla^{2}\psi^{(2)}-9\frac{a^{\prime}}{a}\psi^{(2)^{\prime}}-3\psi^{(2)^{\prime\prime}}\right.$
$\left.-3\frac{a^{\prime}}{a}\nabla^{2}w^{(2)}-\nabla^{2}w^{(2)^{\prime}}+\frac{1}{2}\partial_{k}\partial_{m}D^{km}h^{(2)}+24\frac{a^{\prime\prime}}{a}(\phi^{(1)})^{2}+2\partial_{k}\phi^{(1)}\partial^{k}\phi^{(1)}\right.$
$\left.+4\phi^{(1)}\nabla^{2}\phi^{(1)}+24\frac{a^{\prime}}{a}\phi^{(1)}\phi^{(1)^{\prime}}+6\phi^{(1)^{\prime}}\psi^{(1)^{\prime}}+36\frac{a^{\prime}}{a}\phi^{(1)}\psi^{(1)^{\prime}}+2\partial_{k}\psi^{(1)}\partial^{k}\phi^{(1)}\right.$
$\left.-4\psi^{(1)}\nabla^{2}\phi^{(1)}+12\phi^{(1)}\psi^{(1)^{\prime\prime}}-12\psi^{(1)}\psi^{(1)^{\prime\prime}}-36\frac{a^{\prime\prime}}{a}\psi^{(1)^{\prime}}\psi^{(1)}+6\partial^{k}\psi^{(1)}\partial_{k}\psi^{(1)}\right.$
$\left.+16\psi^{(1)}\nabla^{2}\psi^{(1)}+6\frac{a^{\prime}}{a}\partial_{k}\phi^{(1)}\partial^{k}w^{(1)}+12\frac{a^{\prime}}{a}\phi^{(1)}\nabla^{2}w^{(1)}+4\phi^{(1)}\nabla^{2}w^{(1)^{\prime}}\right.$
$\left.+2\phi^{(1)^{\prime}}\nabla^{2}w^{(1)}+8\partial^{k}\psi^{(1)^{\prime}}\partial_{k}w^{(1)}+2\partial^{k}\psi^{(1)}\partial_{k}w^{(1)^{\prime}}-4\psi^{(1)}\nabla^{2}w^{(1)^{\prime}}\right.$
$\left.-12\frac{a^{\prime}}{a}\psi^{(1)}\nabla^{2}w^{(1)}+4\psi^{(1)^{\prime}}\nabla^{2}w^{(1)}+6\frac{a^{\prime}}{a}\partial^{k}\psi^{(1)}\partial_{k}w^{(1)}-6\frac{a^{\prime\prime}}{a}\partial_{k}w^{(1)}\partial^{k}w^{(1)}\right.$
$\left.-6\frac{a^{\prime}}{a}\partial_{k}w^{(1)}\partial^{k}w^{(1)^{\prime}}+\nabla^{2}w^{(1)}\nabla^{2}w^{(1)}-\partial_{k}\partial_{m}w^{(1)}\partial^{k}\partial^{m}w^{(1)}+2\partial_{k}\phi^{(1)}\partial_{m}D^{mk}h^{(1)}\right.$
$\left.+2\partial_{m}\partial_{k}\phi^{(1)}D^{km}h^{(1)}+4\psi^{(1)}\partial_{k}\partial_{m}D^{mk}h^{(1)}+2\partial_{m}\partial_{k}\psi^{(1)}D^{km}h^{(1)}\right.$
$\left.+6\frac{a^{\prime}}{a}\partial^{k}w^{(1)}\partial_{m}D^{m}_{k}h^{(1)}+6\frac{a^{\prime}}{a}\partial_{k}\partial_{m}w^{(1)}D^{km}h^{(1)}+2\partial_{k}w^{(1)}\partial^{m}D^{k}_{m}h^{(1)^{\prime}}\right.$
$\left.+2\partial_{m}w^{(1)^{\prime}}\partial_{k}D^{km}h^{(1)}+2\partial_{k}\partial_{m}w^{(1)^{\prime}}D^{km}h^{(1)}+\partial_{k}\partial_{m}w^{(1)}D^{km}h^{(1)^{\prime}}\right.$
$\left.-2\partial_{k}\partial^{m}D_{nm}h^{(1)}D^{kn}h^{(1)}-\partial_{k}D^{km}h^{(1)}\partial^{n}D_{mn}h^{(1)}+\nabla^{2}D_{km}h^{(1)}D^{mn}h^{(1)}\right.$
$\left.+\frac{3}{4}\partial^{n}D^{km}h^{(1)}\partial_{n}D_{km}h^{(1)}-\frac{1}{2}\partial_{k}D_{mn}h^{(1)}\partial^{n}D^{mk}h^{(1)}-3\frac{a^{\prime}}{a}D^{km}h^{(1)}D_{km}h^{(1)^{\prime}}\right.$
$\left.-\frac{3}{4}D^{km}h^{(1)^{\prime}}D_{km}h^{(1)^{\prime}}-D^{km}h^{(1)}D_{km}h^{(1)^{\prime\prime}}\right).$ (80)
Appendix C Einstein tensor
$G^{\lambda}_{\mu}=g^{\nu\lambda}\left(R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R\right)$ (81)
C.1 Background components
$G^{(0)0}_{0}$ $=-\frac{3}{a^{2}}\left(\frac{a^{\prime}}{a}\right)^{2},$ (82)
$G^{(0)0}_{i}$ $=G^{(0)i}_{0}=0,$ (83)
$G^{(0)i}_{j}$ $=-\frac{1}{a^{2}}\left[2\frac{a^{\prime\prime}}{a}-\left(\frac{a^{\prime}}{a}\right)^{2}\right]\delta^{i}_{j}.$ (84)
C.2 First-order components
$G^{(1)0}_{0}$ $=\frac{1}{a^{2}}\left[6\left(\frac{a^{\prime}}{a}\right)^{2}\phi^{(1)}+6\frac{a^{\prime}}{a}\psi^{(1)^{\prime}}-2\nabla^{2}\psi^{(1)}+2\frac{a^{\prime}}{a}\nabla^{2}w^{(1)}-\frac{1}{2}\partial^{k}\partial_{m}D^{m}_{k}h^{(1)}\right],$ (85)
$G^{(1)0}_{i}$ $=\frac{1}{a^{2}}\left[-2\frac{a^{\prime}}{a}\partial_{i}\phi^{(1)}-2\partial_{i}\psi^{(1)^{\prime}}-\frac{1}{2}\partial_{k}D^{k}_{i}h^{(1)^{\prime}}\right],$ (86)
$G^{(1)i}_{0}$ $=\frac{1}{a^{2}}\left[4\left(\frac{a^{\prime}}{a}\right)^{2}\partial^{i}w^{(1)}-2\frac{a^{\prime\prime}}{a}\partial^{i}w^{(1)}+2\partial^{i}\psi^{(1)^{\prime}}+2\frac{a^{\prime}}{a}\partial^{i}\phi^{(1)}+\frac{1}{2}\partial_{k}D^{ki}h^{(1)^{\prime}}\right],$ (87)
$G^{(1)i}_{j}$ $=\frac{1}{a^{2}}\left[\left(\left[4\frac{a^{\prime\prime}}{a}-2\left(\frac{a^{\prime}}{a}\right)^{2}\right]\phi^{(1)}+2\frac{a^{\prime}}{a}\phi^{(1)^{\prime}}+\nabla^{2}\phi^{(1)}+4\frac{a^{\prime}}{a}\psi^{(1)^{\prime}}\right.\right.$
$\left.\left.+2\psi^{(1)^{\prime\prime}}-\nabla^{2}\psi^{(1)}+2\frac{a^{\prime}}{a}\nabla^{2}w^{(1)}+\nabla^{2}w^{(1)^{\prime}}-\frac{1}{2}\partial_{k}\partial^{m}D^{k}_{m}h^{(1)}\right)\delta^{i}_{j}\right.$
$\left.-\partial^{i}\partial_{j}\phi^{(1)}+\partial^{i}\partial_{j}\psi^{(1)}-2\frac{a^{\prime}}{a}\partial^{i}\partial_{j}w^{(1)}-\partial^{i}\partial_{j}w^{(1)^{\prime}}+\frac{a^{\prime}}{a}D^{i}_{j}h^{(1)^{\prime}}\right.$
$\left.+\frac{1}{2}D^{i}_{j}h^{(1)^{\prime\prime}}+\frac{1}{2}\partial^{i}\partial_{k}D^{k}_{j}h^{(1)}+\frac{1}{2}\partial_{k}\partial_{j}D^{ik}h^{(1)}-\frac{1}{2}\partial^{k}\partial_{k}D^{i}_{j}h^{(1)}\right].$ (88)
C.3 Second-order components
$G^{(2)0}_{0}$ $=\frac{1}{a^{2}}\left[3\left(\frac{a^{\prime}}{a}\right)^{2}\phi^{(2)}-\nabla^{2}\psi^{(2)}+3\frac{a^{\prime}}{a}\psi^{(2)^{\prime}}+\frac{a^{\prime}}{a}\nabla^{2}w^{(2)}-\frac{1}{4}\partial_{k}\partial_{m}D^{km}h^{(2)}\right.$
$\left.-12\left(\frac{a^{\prime}}{a}\right)^{2}(\phi^{(1)})^{2}+3\left(\frac{a^{\prime}}{a}\right)^{2}\partial_{k}w^{(1)}\partial^{k}w^{(1)}-12\frac{a^{\prime}}{a}\phi^{(1)}\psi^{(1)^{\prime}}-3\partial^{k}\psi^{(1)}\partial_{k}\psi^{(1)}\right.$
$\left.-8\psi^{(1)}\nabla^{2}\psi^{(1)}+12\frac{a^{\prime}}{a}\psi^{(1)}\psi^{(1)^{\prime}}-3(\psi^{(1)})^{2}-4\frac{a^{\prime}}{a}\phi^{(1)}\nabla^{2}w^{(1)}\right.$
$\left.-2\frac{a^{\prime}}{a}\partial^{k}\phi^{(1)}\partial_{k}w^{(1)}-2\frac{a^{\prime}}{a}\partial_{k}\psi^{(1)}\partial^{k}w^{(1)}+4\frac{a^{\prime}}{a}\psi^{(1)}\nabla^{2}w^{(1)}-2\partial^{k}\psi^{(1)^{\prime}}\partial_{k}w^{(1)}\right.$
$\left.-2\psi^{(1)^{\prime}}\nabla^{2}w^{(1)}+\frac{1}{2}\partial_{k}\partial_{m}w^{(1)}\partial^{k}\partial^{m}w^{(1)}-\frac{1}{2}\nabla^{2}w^{(1)}\nabla^{2}w^{(1)}\right.$
$\left.-2\psi^{(1)}\partial_{k}\partial^{m}D^{k}_{m}h^{(1)}+\partial_{k}\partial_{m}\psi^{(1)}D^{km}h^{(1)}-2\frac{a^{\prime}}{a}\partial_{k}\partial_{m}w^{(1)}D^{km}h^{(1)}\right.$
$\left.-\frac{1}{2}\partial_{k}\partial_{m}w^{(1)}D^{km}h^{(1)^{\prime}}-2\frac{a^{\prime}}{a}\partial_{k}w^{(1)}\partial_{m}D^{km}h^{(1)}-\frac{1}{2}\partial_{k}w^{(1)}\partial_{m}D^{km}h^{(1)^{\prime}}\right.$
$\left.-\frac{1}{2}\nabla^{2}D_{mk}h^{(1)}D^{mk}h^{(1)}+\partial_{n}\partial^{k}D_{mk}h^{(1)}D^{mn}h^{(1)}+\frac{1}{2}\partial_{k}D^{km}h^{(1)}\partial^{n}D_{mn}h^{(1)}\right.$
$\left.-\frac{3}{8}\partial^{n}D^{km}h^{(1)}\partial_{n}D_{km}h^{(1)}+\frac{1}{4}\partial_{k}D_{mn}h^{(1)}\partial^{m}D^{kn}h^{(1)}\right.$
$\left.+\frac{a^{\prime}}{a}D^{kn}h^{(1)}D_{kn}h^{(1)^{\prime}}+\frac{1}{8}D^{kn}h^{(1)^{\prime}}D_{kn}h^{(1)^{\prime}}\right],$ (89)
$G^{(2)0}_{i}$ $=\frac{1}{a^{2}}\left[-\frac{a^{\prime}}{a}\partial_{i}\phi^{(2)}-\partial_{i}\psi^{(2)^{\prime}}+\frac{1}{4}\nabla^{2}w_{i}^{(2)}-\frac{1}{4}\partial_{k}D^{k}_{i}h^{(2)^{\prime}}-\frac{1}{4}\nabla^{2}h_{i}^{(2)^{\prime}}\right.$
$\left.8\frac{a^{\prime}}{a}\phi^{(1)}\partial_{i}\phi^{(1)}+4\phi^{(1)}\partial_{i}\psi^{(1)^{\prime}}+2\partial_{i}\phi^{(1)}\psi^{(1)^{\prime}}-4\psi^{(1)^{\prime}}\partial_{i}\psi^{(1)}-4\psi^{(1)^{\prime}}\partial_{i}\psi^{(1)^{\prime}}\right.$
$\left.+\partial_{i}\phi^{(1)}\nabla^{2}w^{(1)}-\partial^{k}\phi^{(1)}\partial_{i}\partial_{k}w^{(1)}+\nabla^{2}\psi^{(1)}\partial_{i}w^{(1)}\right.$
$\left.+\partial_{i}\partial_{k}\psi^{(1)}\partial^{k}w^{(1)}-2\frac{a^{\prime}}{a}\partial_{k}\partial_{i}w^{(1)}\partial^{k}w^{(1)}+\phi^{(1)}\partial^{k}D_{ki}h^{(1)^{\prime}}\right.$
$\left.+\frac{1}{2}\partial^{k}\phi^{(1)}D_{ik}h^{(1)^{\prime}}-\psi^{(1)}\partial_{k}D^{k}_{i}h^{(1)^{\prime}}+\frac{1}{2}\partial_{k}\psi^{(1)}D^{k}_{i}h^{(1)^{\prime}}\right.$
$\left.-\psi^{(1)^{\prime}}\partial_{k}D^{k}_{i}h^{(1)}-\partial_{k}\psi^{(1)^{\prime}}D^{k}_{i}h^{(1)}+\frac{1}{2}\partial_{k}w^{(1)}\partial^{k}\partial^{m}D_{im}h^{(1)}\right.$
$\left.-\frac{1}{2}\partial^{m}w^{(1)}\nabla^{2}D_{im}h^{(1)}+\frac{1}{2}\partial_{k}D^{km}h^{(1)}D_{im}h^{(1)^{\prime}}+\frac{1}{2}D^{km}h^{(1)}\partial_{k}D_{im}h^{(1)^{\prime}}\right.$
$\left.-\frac{1}{4}\partial_{i}D_{mk}h^{(1)}D^{km}h^{(1)^{\prime}}-\frac{1}{2}D^{km}h^{(1)}\partial_{i}D_{mk}h^{(1)^{\prime}}\right],$ (90)
$G^{(2)i}_{0}$ $=\frac{1}{a^{2}}\left[\frac{a^{\prime}}{a}\partial^{i}\phi^{(2)}+\partial^{i}\psi^{(2)^{\prime}}-\frac{1}{4}\nabla^{2}w^{i(2)}+\left(2\left(\frac{a^{\prime}}{a}\right)^{2}-\frac{a^{\prime\prime}}{a}\right)\partial^{i}w^{(2)}\right.$
$\left.+\left(2\left(\frac{a^{\prime}}{a}\right)^{2}-\frac{a^{\prime\prime}}{a}\right)w^{i(2)}+\frac{1}{4}\partial_{k}D^{ki}h^{(2)^{\prime}}+\frac{1}{4}\nabla^{2}h^{i(2)^{\prime}}\right.$
$\left.-4\frac{a^{\prime}}{a}\phi^{(1)}\partial^{i}\phi^{(1)}+4\frac{a^{\prime}}{a}\partial^{i}\phi^{(1)}\psi^{(1)}-2\partial^{i}\phi^{(1)}\psi^{(1)^{\prime}}+4\psi^{(1)^{\prime}}\partial^{i}\psi^{(1)}+8\partial^{i}\psi^{(1)^{\prime}}\psi^{(1)}\right.$
$\left.-\partial^{i}\phi^{(1)}\nabla^{2}w^{(1)}-\partial^{i}\partial_{k}\phi^{(1)}\partial^{k}w^{(1)}+\nabla^{2}\phi^{(1)}\partial^{i}w^{(1)}+\partial^{k}\phi^{(1)}\partial^{i}\partial_{k}w^{(1)}\right.$
$\left.+\left(4\frac{a^{\prime\prime}}{a}-8\left(\frac{a^{\prime}}{a}\right)^{2}\right)\phi^{(1)}\partial^{i}w^{(1)}+2\frac{a^{\prime}}{a}\phi^{(1)^{\prime}}\partial^{i}w^{(1)}+2\phi^{(1)^{\prime\prime}}\partial^{i}w^{(1)}\right.$
$\left.+\left(8\left(\frac{a^{\prime}}{a}\right)^{2}-4\frac{a^{\prime\prime}}{a}\right)\psi^{(1)}\partial^{i}w^{(1)}-2\frac{a^{\prime}}{a}\psi^{(1)^{\prime}}\partial^{i}w^{(1)}+\nabla^{2}w^{(1)^{\prime}}\partial^{i}w^{(1)}\right.$
$\left.-\partial^{i}\partial_{k}w^{(1)^{\prime}}\partial^{k}w^{(1)}-\frac{1}{2}\partial^{k}\phi^{(1)}D^{i}_{k}h^{(1)^{\prime}}-2\frac{a^{\prime}}{a}\partial_{k}\phi^{(1)}D^{ki}h^{(1)}-\frac{1}{2}\partial_{k}\psi^{(1)}D^{ki}h^{(1)^{\prime}}\right.$
$\left.+2\psi^{(1)}\partial_{k}D^{ki}h^{(1)^{\prime}}+\psi^{(1)^{\prime}}\partial_{k}D^{ki}h^{(1)}-\partial_{k}\psi^{(1)^{\prime}}D^{ki}h^{(1)}\right.$
$\left.+\left(2\frac{a^{\prime\prime}}{a}-4\left(\frac{a^{\prime}}{a}\right)^{2}\right)\partial_{k}w^{(1)}D^{ik}h^{(1)}+\frac{a^{\prime}}{a}\partial^{k}w^{(1)}D^{i}_{k}h^{(1)^{\prime}}+\frac{1}{2}\partial^{k}w^{(1)}D^{i}_{k}h^{(1)^{\prime\prime}}\right.$
$\left.-\frac{1}{2}\partial_{k}D^{km}h^{(1)}D^{i}_{m}h^{(1)^{\prime}}-\frac{1}{2}D^{km}h^{(1)}\partial_{k}D^{i}_{m}h^{(1)^{\prime}}+\frac{1}{4}\partial^{i}D_{mk}h^{(1)}D^{km}h^{(1)^{\prime}}\right.$
$\left.+\frac{1}{2}D^{km}h^{(1)}\partial^{i}D_{mk}h^{(1)^{\prime}}-\frac{1}{2}D^{ik}h^{(1)}\partial_{m}D^{m}_{k}h^{(1)^{\prime}}\right],$ (91)
$G^{d,i(2)}_{j}$ $=a^{-2}\left[\frac{1}{2}\nabla^{2}\phi^{(2)}+\left[2\frac{a^{\prime\prime}}{a}-\left(\frac{a^{\prime}}{a}\right)^{2}\right]\phi^{(2)}+\frac{a^{\prime}}{a}\phi^{(2)^{\prime}}-\frac{1}{2}\nabla^{2}\psi^{(2)}+\psi^{(2)^{\prime\prime}}\right.$
$\left.+2\frac{a^{\prime}}{a}\psi^{(2)}+\frac{a^{\prime}}{a}\nabla^{2}w^{(2)}+\frac{1}{2}\nabla^{2}w^{(2)^{\prime}}-\frac{1}{4}\partial_{k}\partial_{m}D^{km}h^{(2)}\right.$
$\left.+\left[4\left(\frac{a^{\prime}}{a}\right)^{2}-8\frac{a^{\prime\prime}}{a}\right](\phi^{(1)})^{2}-8\frac{a^{\prime}}{a}\phi^{(1)}\phi^{(1)^{\prime}}-\partial_{k}\phi^{(1)}\partial^{k}\phi^{(1)}-2\phi^{(1)}\nabla^{2}\phi^{(1)}\right.$
$\left.-4\phi^{(1)}\psi^{(1)^{\prime\prime}}-2\phi^{(1)^{\prime}}\psi^{(1)^{\prime}}-8\frac{a^{\prime}}{a}\phi^{(1)}\psi^{(1)^{\prime}}-2\partial_{k}\psi^{(1)}\partial^{k}\psi^{(1)}\right.$
$\left.-4\psi^{(1)}\nabla^{2}\psi^{(1)}+(\psi^{(1)^{\prime}})^{2}+8\frac{a^{\prime}}{a}\psi^{(1)}\psi^{(1)^{\prime}}+4\psi^{(1)}\psi^{(1)^{\prime\prime}}\right.$
$\left.+2\psi^{(1)}\nabla^{2}\phi^{(1)}-\phi^{(1)^{\prime}}\nabla^{2}w^{(1)}-2\phi^{(1)}\nabla^{2}w^{(1)^{\prime}}-2\frac{a^{\prime}}{a}\partial_{k}w^{(1)}\partial^{k}w^{(1)^{\prime}}\right.$
$\left.-\frac{1}{2}\nabla^{2}w^{(1)}\nabla^{2}w^{(1)}+\frac{1}{2}\partial^{m}\partial^{k}w^{(1)}\partial_{m}\partial_{k}w^{(1)}+4\frac{a^{\prime}}{a}\psi^{(1)}\nabla^{2}w^{(1)}+2\psi^{(1)}\nabla^{2}w^{(1)^{\prime}}\right.$
$\left.-2\partial^{k}\psi^{(1)^{\prime}}\partial_{k}w^{(1)}-\psi^{(1)^{\prime}}\nabla^{2}w^{(1)}-\partial_{k}\partial_{m}\phi^{(1)}D^{km}h^{(1)}-\partial_{k}\phi^{(1)}\partial_{m}D^{km}h^{(1)}\right.$
$\left.-\partial_{k}\psi^{(1)}\partial_{m}D^{km}h^{(1)}-\partial_{k}w^{(1)}\partial^{m}D^{k}_{m}h^{(1)^{\prime}}-\partial_{k}w^{(1)^{\prime}}\partial_{m}D^{km}h^{(1)}\right.$
$\left.+\partial_{k}\partial_{m}w^{(1)^{\prime}}D^{km}h^{(1)}-2\frac{a^{\prime}}{a}\partial^{k}w^{(1)}\partial_{m}D^{m}_{k}h^{(1)^{\prime}}-2\frac{a^{\prime}}{a}\partial^{k}\partial_{m}w^{(1)}D^{m}_{k}h^{(1)}\right.$
$\left.-\frac{1}{2}\partial_{k}\partial_{m}w^{(1)}D^{km}h^{(1)^{\prime}}-2\psi^{(1)}\partial_{k}\partial_{m}D^{km}h^{(1)}+\partial_{k}\partial^{n}D_{mn}h^{(1)}D^{km}h^{(1)}\right.$
$\left.-\frac{1}{2}\nabla^{2}D_{mn}h^{(1)}D^{mn}h^{(1)}+\frac{1}{2}\partial_{k}D^{k}_{m}h^{(1)}\partial^{n}D_{n}^{m}h^{(1)}+\frac{1}{2}D^{mk}h^{(1)}D_{mk}h^{(1)^{\prime\prime}}\right.$
$\left.-\frac{3}{8}\partial^{n}D_{km}h^{(1)}\partial^{n}D^{km}h^{(1)}+\frac{3}{8}D^{mk}h^{(1)^{\prime}}D_{mk}h^{(1)^{\prime}}\right.$
$\left.+\frac{a^{\prime}}{a}D^{mk}h^{(1)}D_{mk}h^{(1)^{\prime}}+\frac{1}{4}\partial^{n}D^{km}h^{(1)}\partial_{m}D_{kn}h^{(1)}\right]\delta^{i}_{j},$
$G^{nd,i(2)}_{j}$ $=a^{-2}\left[-\frac{1}{2}\partial^{i}\partial_{j}\phi^{(2)}+\frac{1}{2}\partial^{i}\partial_{j}\psi^{(2)}-\frac{a^{\prime}}{a}\partial^{i}\partial_{j}w^{(2)}-\frac{1}{2}\partial^{i}\partial_{j}w^{(2)^{\prime}}\right.$
$\left.-\frac{1}{2}\frac{a^{\prime}}{a}\left(\partial^{i}w_{j}^{(2)}+\partial_{j}w^{i(2)}\right)-\frac{1}{4}\left(\partial^{i}w_{j}^{(2)^{\prime}}+\partial_{j}w^{i(2)^{\prime}}\right)\right.$
$\left.+\frac{1}{2}\frac{a^{\prime}}{a}\left(D^{i}_{j}h^{(2)^{\prime}}+\partial^{i}h_{j}^{(2)^{\prime}}+\partial_{j}h^{i(2)^{\prime}}+h^{i(2)^{\prime}}_{j}\right)+\frac{1}{2}\partial_{k}\partial^{i}D^{k}_{j}h^{(2)}\right.$
$\left.-\frac{1}{4}\nabla^{2}D^{i}_{j}h^{(2)}-\frac{1}{4}\nabla^{2}h^{i(2)}_{j}\right.$
$\left.+\frac{1}{4}\left(D^{i}_{j}h^{(2)^{\prime\prime}}++\partial^{i}h_{j}^{(2)^{\prime\prime}}+\partial_{j}h^{i(2)^{\prime\prime}}+h^{i(2)^{\prime\prime}}_{j}\right)\right.$
$\left.+\partial^{i}\phi^{(1)}\partial_{j}\phi^{(1)}+2\phi^{(1)}\partial^{i}\partial_{j}\phi^{(1)}-2\psi^{(1)}\partial^{i}\partial_{j}\phi^{(1)}-\partial_{j}\phi^{(1)}\partial^{i}\psi^{(1)}\right.$
$\left.-\partial^{i}\phi^{(1)}\partial_{j}\psi^{(1)}+3\partial^{i}\psi^{(1)}\partial_{j}\psi^{(1)}+4\psi^{(1)}\partial^{i}\partial_{j}\psi^{(1)}+2\frac{a^{\prime}}{a}\partial^{i}w^{(1)}\partial_{j}{(\phi^{(1)}}\right.$
$\left.+4\frac{a^{\prime}}{a}\phi^{(1)}\partial^{j}\partial_{i}w^{(1)}+\phi^{(1)^{\prime}}\partial^{i}\partial_{j}w^{(1)}+2\phi^{(1)}\partial^{i}\partial_{j}w^{(1)^{\prime}}+\nabla^{2}w^{(1)}\partial^{i}\partial_{j}w^{(1)}\right.$
$\left.-\partial_{j}\partial^{k}w^{(1)}\partial^{i}\partial_{k}w^{(1)}-2\frac{a^{\prime}}{a}\partial^{i}\psi^{(1)}\partial_{j}w^{(1)}-2\frac{a^{\prime}}{a}\partial^{i}w^{(1)}\partial_{j}\psi^{(1)}-\partial^{i}\psi^{(1)^{\prime}}\partial_{j}w^{(1)}\right.$
$\left.+\frac{a^{\prime}}{a}\partial^{i}w^{(1)}\partial_{j}\psi^{(1)^{\prime}}-\partial^{i}\psi^{(1)}\partial_{j}w^{(1)^{\prime}}-\frac{a^{\prime}}{a}\partial^{i}w^{(1)^{\prime}}\partial_{j}\psi^{(1)}-2\psi^{(1)}\partial^{i}\partial_{j}w^{(1)^{\prime}}\right.$
$\left.+\psi^{(1)^{\prime}}\partial^{i}\partial_{j}w^{(1)}-2\frac{a^{\prime}}{a}\psi^{(1)}\partial^{i}\partial_{j}w^{(1)}-2\frac{a^{\prime}}{a}\phi^{(1)}D^{i}_{j}h^{(1)^{\prime}}-\frac{1}{2}\phi^{(1)^{\prime}}D^{i}_{j}h^{(1)^{\prime}}\right.$
$\left.-\phi^{(1)}D^{i}_{j}h^{(1)^{\prime\prime}}+\frac{1}{2}\partial_{k}\phi^{(1)}\partial^{i}D^{k}_{j}h^{(1)}+\frac{1}{2}\partial_{k}\phi^{(1)}\partial_{j}D^{ki}h^{(1)}-\frac{1}{2}\partial_{k}\phi^{(1)}\partial^{k}D^{i}_{j}h^{(1)}\right.$
$\left.+\partial_{j}\partial_{k}\phi^{(1)}D^{ki}h^{(1)}+\frac{1}{2}\psi^{(1)^{\prime}}D^{i}_{j}h^{(1)^{\prime}}+\psi^{(1)^{\prime\prime}}D^{i}_{j}h^{(1)}+2\frac{a^{\prime}}{a}\psi^{(1)^{\prime}}D^{i}_{j}h^{(1)}\right.$
$\left.+\frac{1}{2}\partial_{k}\psi^{(1)}\partial^{i}D^{k}_{j}h^{(1)}+2\frac{a^{\prime}}{a}\psi^{(1)}D^{i}_{j}h^{(1)^{\prime}}+\psi^{(1)}D^{i}_{j}h^{(1)^{\prime\prime}}+\frac{1}{2}\partial_{k}\psi^{(1)}\partial_{j}D^{ki}h^{(1)}\right.$
$\left.-\frac{3}{2}\partial_{k}\psi^{(1)}\partial^{k}D^{i}_{j}h^{(1)}+2\psi^{(1)}\partial_{k}\partial^{i}D^{k}_{j}h^{(1)}+2\psi^{(1)}\partial_{k}\partial_{j}D^{ki}h^{(1)}\right.$
$\left.-2\psi^{(1)}\nabla^{2}D^{i}_{j}h^{(1)}-\nabla^{2}\psi^{(1)}D^{i}_{j}h^{(1)}+\partial^{i}\psi^{(1)}\partial_{k}D^{k}_{j}h^{(1)}\right.$
$\left.+\partial_{j}\psi^{(1)}\partial_{k}D^{ki}h^{(1)}+\partial_{k}\partial^{i}\psi^{(1)}D^{k}_{j}h^{(1)}+\frac{1}{2}\partial^{i}w^{(1)}\partial_{k}D^{k}_{j}h^{(1)^{\prime}}\right.$
$\left.+\frac{1}{2}\partial_{k}\partial^{i}w^{(1)}D^{k}_{j}h^{(1)^{\prime}}+\frac{1}{2}\partial_{k}\partial_{j}w^{(1)}D^{ki}h^{(1)^{\prime}}-\frac{1}{2}\nabla^{2}w^{(1)}D^{i}_{j}h^{(1)^{\prime}}\right.$
$\left.+\frac{1}{2}\partial^{k}w^{(1)}\partial^{i}D_{kj}h^{(1)^{\prime}}+\frac{1}{2}\partial^{k}w^{(1)}\partial_{j}D^{i}_{k}h^{(1)^{\prime}}\right.$
$\left.-\partial^{k}w^{(1)}\partial_{k}D^{i}_{j}h^{(1)^{\prime}}+\frac{1}{2}\partial^{k}w^{(1)^{\prime}}\partial^{i}D_{kj}h^{(1)}+\frac{1}{2}\partial^{k}w^{(1)^{\prime}}\partial_{j}D^{i}_{k}h^{(1)}\right.$
$\left.-\frac{1}{2}\partial^{k}w^{(1)^{\prime}}\partial_{k}D^{i}_{j}h^{(1)}+\partial_{k}\partial_{j}w^{(1)^{\prime}}D^{ik}h^{(1)}+\frac{a^{\prime}}{a}\partial^{k}w^{(1)}\partial^{i}D_{kj}h^{(1)}\right.$
$\left.+\frac{a^{\prime}}{a}\partial^{k}w^{(1)}\partial_{j}D^{i}_{k}h^{(1)}-\frac{a^{\prime}}{a}\partial^{k}w^{(1)}\partial_{k}D^{i}_{j}h^{(1)}+2\frac{a^{\prime}}{a}\partial_{k}\partial_{j}w^{(1)}D^{ik}h^{(1)}\right.$
$\left.-\frac{1}{2}D^{ki}h^{(1)^{\prime}}D_{kj}h^{(1)^{\prime}}-\frac{1}{2}\partial^{i}D_{mj}h^{(1)}\partial_{k}D^{km}h^{(1)}-\frac{1}{2}\partial_{j}D^{i}_{m}h^{(1)}\partial_{k}D^{km}h^{(1)}\right.$
$\left.+\frac{1}{2}\partial_{m}D^{i}_{j}h^{(1)}\partial_{k}D^{km}h^{(1)}-\frac{1}{2}\partial_{k}\partial^{i}D_{mj}h^{(1)}D^{km}h^{(1)}\right.$
$\left.-\frac{1}{2}\partial_{k}\partial_{j}D^{i}_{m}h^{(1)}D^{km}h^{(1)}+\frac{1}{2}\partial_{k}\partial_{m}D^{i}_{j}h^{(1)}D^{km}h^{(1)}\right.$
$\left.+\frac{1}{2}D^{km}h^{(1)}\partial^{i}\partial_{j}D_{km}h^{(1)}+\frac{1}{4}\partial^{i}D^{mk}h^{(1)}\partial_{j}D_{mk}h^{(1)}\right.$
$\left.-\frac{a^{\prime}}{a}D_{kj}h^{(1)^{\prime}}D^{ik}h^{(1)}-\frac{1}{2}D_{kj}h^{(1)^{\prime\prime}}D^{ki}h^{(1)}\right.$
$\left.-\partial_{m}\partial_{k}D^{m}_{j}h^{(1)}D^{ki}h^{(1)}+\frac{1}{2}\partial_{m}\partial^{m}D_{kj}h^{(1)}D^{ki}h^{(1)}\right.$
$\left.+\frac{1}{2}\partial_{m}D^{ik}h^{(1)}\partial^{m}D_{kj}h^{(1)}-\frac{1}{2}\partial_{m}D^{ik}h^{(1)}\partial_{k}D^{m}_{j}h^{(1)}\right].$ (92)
Appendix D Energy-momentum tensor
The energy-momentum tensor describes the perturbed matter quantities. The following is the expression for a fluid,
$T^{\mu}_{\nu}=(\rho+p)u^{\mu}u_{\nu}+p\delta^{\mu}_{\nu}+\pi^{\mu}_{\nu}$ (93)
where the perturbed energy density, $\rho$, and pressure, $p$.
$\rho=\bar{\rho}+\sum_{r=1}^{\infty}\frac{1}{r!}\delta^{(r)}\rho,$ (94)
$P=\bar{P}+\sum_{r=1}^{\infty}\frac{1}{r!}\delta^{(r)}P.$ (95)
The 4 - velocity, $u^{\mu}$, can be defined in the following way,
$u^{\mu}=\frac{1}{a}\left(\delta^{\mu}_{0}+\sum_{r=1}^{\infty}\frac{1}{r!}v^{\mu}_{(r)}\right),$ (96)
$u^{0}=\frac{1}{a}\left(1+v^{(1)0}+\frac{1}{2}v^{(2)0}\right),$ (97)
$u^{i}=\frac{1}{a}\left(v^{(1)i}+\frac{1}{2}v^{(2)i}\right),$ (98)
where $u^{\mu}$ is subject to normalization condition $u^{\mu}u^{\nu}g_{\mu\nu}=-1$. Using this condition, the components of 4-velocity, $u^{0}$, $u^{i}$, $u_{0}$ and $u_{i}$, up to second order can be found. First by finding, $v^{(1)0}$ and $v^{(2)0}$. For first and second order perturbation,
$u^{0}u^{0}g_{00}$ $=-1$
$\left(1+v^{(1)0}\right)^{2}\left(-1-2\phi^{(1)}\right)$ $=-1$
$v^{(1)0}$ $=-\phi^{(1)}$ (99)
$u^{0}u^{0}g_{00}+2u^{0}u^{i}g_{0i}+u^{i}u^{j}g_{ij}=-1$
$\left(1+v^{(1)0}+\frac{1}{2}v^{(2)0}\right)^{2}\left(-1-2\phi^{(1)}-\phi^{(2)}\right)$
$+2\left(1+v^{(1)0}+\frac{1}{2}v^{(2)0}\right)\left(v^{(1)i}+\frac{1}{2}v^{(2)i}\right)\left(\partial_{i}w^{(1)}+\frac{1}{2}\partial_{i}w^{(2)}+\frac{1}{2}w_{i}^{(2)}\right)$
$+\left(v^{(1)i}+\frac{1}{2}v^{(2)i}\right)\left(v^{(1)j}+\frac{1}{2}v^{(2)j}\right)\left[\left(1-2\psi^{(1)}-\psi^{(2)}\right)\delta_{ij}\right.$
$\left.+D_{ij}\left(h^{(1)}+\frac{1}{2}h^{(2)}\right)+\frac{1}{2}\left(\partial_{i}h_{(2)j}+\partial_{j}h_{(2)i}+h^{(2)}_{ij}\right)\right]=-1,$
substituting in the expression for $v^{(1)0}$ we get,
$v^{0(2)}=-\phi^{(2)}+3\left(\phi^{(1)}\right)^{2}+v^{(1)i}v_{i(1)}+2v^{(1)i}\partial_{i}w^{(1)}.$ (100)
By using the condition $u^{\mu}u_{\mu}=-1$, all components of the 4-velocity is found to be the following,
$u^{0}$ $=a^{-1}\left[1-\phi^{(1)}-\frac{1}{2}\phi^{(2)}+\frac{3}{2}\left(\phi^{(1)}\right)^{2}+\frac{1}{2}v^{(1)k}v_{(1)k}+v_{k}^{(1)}\partial^{k}w^{(1)}\right],$
$u^{i}$ $=a^{-1}\left[v^{i(1)}+\frac{1}{2}v^{i(2)}\right],$ (101)
$u_{0}$ $=a\left[-1-\phi^{(1)}-\frac{1}{2}\phi^{(2)}+\frac{1}{2}\left(\phi^{(1)}\right)^{2}-\frac{1}{2}v^{i(1)}v_{i(1)}\right],$
$u_{i}$ $=a\left[\partial_{i}w^{(1)}-v_{i}^{(1)}+\frac{1}{2}\left(\partial_{i}w^{(2)}+w_{i}^{(2)}+v_{i}^{(2)}\right)\right.$
$\left.-\partial_{i}w^{(1)}\phi^{(1)}-2v^{k(1)}\psi^{(1)}+v^{k(1)}D_{ik}h^{(1)}\right],$ (102)
The anisotropic stress tensor, $\pi_{\mu\nu}$, can also be split into first and second order parts,
$\pi_{\mu\nu}=\pi^{(1)}_{\mu\nu}+\frac{1}{2}\pi^{(2)}_{\mu\nu},$ (103)
and it is subject to the constraint $\pi^{\mu}_{\mu}=0$ and $\pi_{\mu\nu}u^{\mu}=0$. The anisotropic stress tensor decomposes into trace-free scalar part,$\Pi$, vector part $\Pi_{i}$, and tensor, $\Pi_{ij}$ part [18].
$\pi_{ij}=a^{2}\left[\Pi_{,ij}-\frac{1}{3}\delta_{ij}\nabla^{2}\Pi+\frac{1}{2}\left(\Pi_{i,j}+\Pi_{j,i}\right)+\Pi_{ij}\right]$ (104)
From the definitions above, the components of the stress energy tensor can be found.
D.1 Time - time component of energy-momentum tensor
$T^{0}_{0}$ $=T^{(0)0}_{0}+T^{(1)0}_{0}+T^{(2)0}_{0},$
$=(\rho+p)u^{0}u_{0}+p\delta^{0}_{0}+\pi^{0}_{0},$
$=\left(\bar{\rho}+\rho^{(1)}+\frac{1}{2}\rho^{(2)}+\bar{P}+P^{(1)}+\frac{1}{2}P^{(2)}\right)\left(-v^{k(1)}v_{k(1)}-\partial^{k}w^{(1)}v_{k}^{(1)}\right)$
$-\left(\bar{\rho}+\rho^{(1)}+\frac{1}{2}\rho^{(2)}\right).$ (105)
$T^{(0)0}_{0}=-\bar{\rho},$ (106)
$T^{(1)0}_{0}=-\rho^{(1)},$ (107)
$T^{0(2)}_{0}=-\rho^{(2)}-2\left(\bar{\rho}+\bar{P}\right)v_{k}^{(1)}\left(v^{k(1)}+\partial^{k}w^{(1)}\right).$ (108)
D.2 Time - space component of energy-momentum tensor
$T^{0}_{i}$ $=T^{(0)0}_{i}+T^{(1)0}_{i}+T^{(2)0}_{i},$
$=(\rho+p)u^{0}u_{i}+p\delta^{0}_{i}+\pi^{0}_{i},$
$=\left(\bar{\rho}+\rho^{(1)}+\frac{1}{2}\rho^{(2)}+\bar{P}+P^{(1)}+\frac{1}{2}P^{(2)}\right)\left(\partial_{i}w^{(1)}+v_{i}^{(1)}\right.$
$\left.+\frac{1}{2}\left(\partial_{i}w^{(2)}+w_{i}^{(2)}+v_{i}^{(2)}\right)-v_{i}^{(1)}\phi^{(1)}-2\partial_{i}w^{(1)}\phi^{(1)}-2v^{k(1)}\psi^{(1)}+v^{k(1)}D_{ik}h^{(1)}\right)$
$+2a^{-2}\pi^{(1)}_{ik}\left(\partial^{k}w^{(1)}+v^{k(1)}\right).$ (109)
$T^{(0)0}_{i}=0,$ (110)
$T^{(1)0}_{i}=\left(\bar{\rho}+\bar{P}\right)\left(\partial_{i}w^{(1)}+v_{i}^{(1)}\right),$ (111)
$T^{(2)0}_{i}$ $=\left(\bar{\rho}+\bar{P}\right)\left[\partial_{i}w^{(2)}+w_{i}^{(2)}+v_{i}^{(2)}-2\phi^{(1)}\left(v_{i}^{(1)}+2\partial_{i}w^{(1)}\right)\right.$
$\left.-4v^{k(1)}\psi^{(1)}\delta_{ik}+2v^{k(1)}D_{ik}h^{(1)}\right]$
$+2\left(\rho^{(1)}+P^{(1)}\right)\left(\partial_{i}w^{(1)}+v_{i}^{(1)}\right)+a^{-2}\pi^{(1)}_{ik}\left(\partial^{k}w^{(1)}+v^{k(1)}\right).$ (112)
D.3 Space - space component of energy-momentum tensor
$T^{i}_{j}$ $=T^{(0)i}_{j}+T^{(1)i}_{j}+T^{(2)i}_{j},$
$=(\rho+p)u^{i}u_{j}+p\delta^{i}_{j}+\pi^{i}_{j},$
$=\left(\bar{\rho}+\rho^{(1)}+\frac{1}{2}\rho^{(2)}+\bar{P}+P^{(1)}+\frac{1}{2}P^{(2)}\right)\left(v^{i(1)}\partial_{j}w^{(1)}+v^{i(1)}v_{j(1)}\right)$
$+\left(\bar{P}+P^{(1)}+\frac{1}{2}P^{(2)}\right)\delta^{i}_{j}+a^{-2}\pi^{(1)i}_{j}$
$+\frac{1}{2}a^{-2}\left[\pi^{(2)i}_{j}+\left(4\psi^{(1)}\delta^{ik}-2D^{ik}h^{(1)}\right)\pi^{(1)}_{jk}\right].$ (113)
$T^{(0)i}_{j}=\bar{P}\delta^{i}_{j},$ (114)
$T^{(1)i}_{j}=P^{(1)}\delta^{i}_{j}+a^{-2}\pi^{(1)i}_{j},$ (115)
$T^{(2)i}_{j}$ $=P^{(2)}\delta^{i}_{j}+2\left(\bar{\rho}+\bar{P}\right)v^{i(1)}\left(\partial_{j}w^{(1)}+v_{j(1)}\right)$
$+a^{-2}\left[\pi^{(2)i}_{j}+\left(4\psi^{(1)}\delta^{ik}-2D^{ik}h^{(1)}\right)\pi^{(1)}_{jk}\right].$ (116)
Appendix E Source term of SIGW
The source term of the SIGW in a generic gauge is given by,
$S^{i}_{j}=$ $\left.\partial^{i}\phi^{(1)}\partial_{j}\phi^{(1)}+2\phi^{(1)}\partial^{i}\partial_{j}\phi^{(1)}-2\psi^{(1)}\partial^{i}\partial_{j}\phi^{(1)}-\partial_{j}\phi^{(1)}\partial^{i}\psi^{(1)}-\partial^{i}\phi^{(1)}\partial_{j}\psi^{(1)}\right.$
$\left.+3\partial^{i}\psi^{(1)}\partial_{j}\psi^{(1)}+4\psi^{(1)}\partial^{i}\partial_{j}\psi^{(1)}+2\mathcal{H}\partial^{i}w^{(1)}\partial_{j}{\phi^{(1)}}+4\mathcal{H}\phi^{(1)}\partial^{j}\partial_{i}w^{(1)}\right.$
$\left.+\phi^{(1)^{\prime}}\partial^{i}\partial_{j}w^{(1)}+2\phi^{(1)}\partial^{i}\partial_{j}w^{(1)^{\prime}}+\nabla^{2}w^{(1)}\partial^{i}\partial_{j}w^{(1)}-\partial_{j}\partial^{k}w^{(1)}\partial^{i}\partial_{k}w^{(1)}\right.$
$\left.-2\mathcal{H}\partial^{i}\psi^{(1)}\partial_{j}w^{(1)}-2\mathcal{H}\partial^{i}w^{(1)}\partial_{j}\psi^{(1)}-\partial^{i}\psi^{(1)^{\prime}}\partial_{j}w^{(1)}+\partial^{i}w^{(1)}\partial_{j}\psi^{(1)^{\prime}}\right.$
$\left.-\partial^{i}\psi^{(1)}\partial_{j}w^{(1)^{\prime}}-\partial^{i}w^{(1)^{\prime}}\partial_{j}\psi^{(1)}-2\psi^{(1)}\partial^{i}\partial_{j}w^{(1)^{\prime}}+\psi^{(1)^{\prime}}\partial^{i}\partial_{j}w^{(1)}\right.$
$\left.-4\mathcal{H}\psi^{(1)}\partial^{i}\partial_{j}w^{(1)}-2\mathcal{H}\phi^{(1)}D^{i}_{j}h^{(1)^{\prime}}-\frac{1}{2}\phi^{(1)^{\prime}}D^{i}_{j}h^{(1)^{\prime}}-\phi^{(1)}D^{i}_{j}h^{(1)^{\prime\prime}}\right.$
$\left.+\frac{1}{2}\partial_{k}\phi^{(1)}\partial^{i}D^{k}_{j}h^{(1)}+\frac{1}{2}\partial_{k}\phi^{(1)}\partial_{j}D^{ki}h^{(1)}-\frac{1}{2}\partial_{k}\phi^{(1)}\partial^{k}D^{i}_{j}h^{(1)}+\partial_{j}\partial_{k}\phi^{(1)}D^{ki}h^{(1)}\right.$
$\left.\frac{1}{2}\psi^{(1)^{\prime}}D^{i}_{j}h^{(1)^{\prime}}+\psi^{(1)^{\prime\prime}}D^{i}_{j}h^{(1)}+2\mathcal{H}\psi^{(1)^{\prime}}D^{i}_{j}h^{(1)}+\frac{1}{2}\partial_{k}\psi^{(1)}\partial^{i}D^{k}_{j}h^{(1)}\right.$
$\left.+2\mathcal{H}\psi^{(1)}D^{i}_{j}h^{(1)^{\prime}}+\psi^{(1)}D^{i}_{j}h^{(1)^{\prime\prime}}+\frac{1}{2}\partial_{k}\psi^{(1)}\partial_{j}D^{ki}h^{(1)}-\frac{3}{2}\partial_{k}\psi^{(1)}\partial^{k}D^{i}_{j}h^{(1)}\right.$
$\left.+2\psi^{(1)}\partial_{k}\partial^{i}D^{k}_{j}h^{(1)}+2\psi^{(1)}\partial_{k}\partial_{j}D^{ki}h^{(1)}-2\psi^{(1)}\nabla^{2}D^{i}_{j}h^{(1)}-\nabla^{2}\psi^{(1)}D^{i}_{j}h^{(1)}\right.$
$\left.+\partial^{i}\psi^{(1)}\partial_{k}D^{k}_{j}h^{(1)}+\partial_{j}\psi^{(1)}\partial_{k}D^{ki}h^{(1)}+\partial_{k}\partial^{i}\psi^{(1)}D^{k}_{j}h^{(1)}+\frac{1}{2}\partial^{i}w^{(1)}\partial_{k}D^{k}_{j}h^{(1)^{\prime}}\right.$
$\left.+\frac{1}{2}\partial_{k}\partial^{i}w^{(1)}D^{k}_{j}h^{(1)^{\prime}}+\frac{1}{2}\partial_{k}\partial_{j}w^{(1)}D^{ki}h^{(1)^{\prime}}-\frac{1}{2}\nabla^{2}w^{(1)}D^{i}_{j}h^{(1)^{\prime}}\right.$
$\left.+\frac{1}{2}\partial^{k}w^{(1)}\partial^{i}D_{kj}h^{(1)^{\prime}}+\frac{1}{2}\partial^{k}w^{(1)}\partial_{j}D^{i}_{k}h^{(1)^{\prime}}-\partial^{k}w^{(1)}\partial_{k}D^{i}_{j}h^{(1)^{\prime}}\right.$
$\left.+\frac{1}{2}\partial^{k}w^{(1)^{\prime}}\partial^{i}D_{kj}h^{(1)}+\frac{1}{2}\partial^{k}w^{(1)^{\prime}}\partial_{j}D^{i}_{k}h^{(1)}-\frac{1}{2}\partial^{k}w^{(1)^{\prime}}\partial_{k}D^{i}_{j}h^{(1)}\right.$
$\left.+\partial_{k}\partial_{j}w^{(1)^{\prime}}D^{ik}h^{(1)}+\mathcal{H}\partial^{k}w^{(1)}\partial^{i}D_{kj}h^{(1)}+\mathcal{H}\partial^{k}w^{(1)}\partial_{j}D^{i}_{k}h^{(1)}\right.$
$\left.-\mathcal{H}\partial^{k}w^{(1)}\partial_{k}D^{i}_{j}h^{(1)}+2\mathcal{H}\partial_{k}\partial_{j}w^{(1)}D^{ik}h^{(1)}-\frac{1}{2}D^{ki}h^{(1)^{\prime}}D_{kj}h^{(1)^{\prime}}\right.$
$\left.-\frac{1}{2}\partial^{i}D_{mj}h^{(1)}\partial_{k}D^{km}h^{(1)}-\frac{1}{2}\partial_{j}D^{i}_{m}h^{(1)}\partial_{k}D^{km}h^{(1)}+\frac{1}{2}\partial_{m}D^{i}_{j}h^{(1)}\partial_{k}D^{km}h^{(1)}\right.$
$\left.-\frac{1}{2}\partial_{k}\partial^{i}D_{mj}h^{(1)}D^{km}h^{(1)}-\frac{1}{2}\partial_{k}\partial_{j}D^{i}_{m}h^{(1)}D^{km}h^{(1)}+\frac{1}{2}\partial_{k}\partial_{m}D^{i}_{j}h^{(1)}D^{km}h^{(1)}\right.$
$\left.+\frac{1}{2}D^{km}h^{(1)}\partial^{i}\partial_{j}D_{km}h^{(1)}+\frac{1}{4}\partial^{i}D^{mk}h^{(1)}\partial_{j}D_{mk}h^{(1)}\right.$
$\left.-\mathcal{H}D_{kj}h^{(1)^{\prime}}D^{ik}h^{(1)}-\frac{1}{2}D_{kj}h^{(1)^{\prime\prime}}D^{ki}h^{(1)}-\partial_{m}\partial_{k}D^{m}_{j}h^{(1)}D^{ki}h^{(1)}\right.$
$\left.+\frac{1}{2}\partial_{m}\partial^{m}D_{kj}h^{(1)}D^{ki}h^{(1)}+\frac{1}{2}\partial_{m}D^{ik}h^{(1)}\partial^{m}D_{kj}h^{(1)}-\frac{1}{2}\partial_{m}D^{ik}h^{(1)}\partial_{k}D^{m}_{j}h^{(1)}\right.$
$-\frac{4}{3\mathcal{H}^{2}\left(1+w\right)}\left[\partial^{i}\left(\psi^{(1)^{\prime}}+\mathcal{H}\phi^{(1)}\right)\partial_{j}\left(\psi^{(1)^{\prime}}+\mathcal{H}\phi^{(1)}\right)\right]$
$-\frac{1}{3\mathcal{H}^{2}\left(1+w\right)}\left[\partial^{i}\left(\psi^{(1)^{\prime}}+\mathcal{H}\phi^{(1)}\right)\partial_{k}D^{k}_{j}h^{(1)^{\prime}}\right.$
$\left.+\partial_{j}\left(\psi^{(1)^{\prime}}+\mathcal{H}\phi^{(1)}\right)\partial_{k}D^{ki}h^{(1)^{\prime}}\right]$
$-\frac{1}{6\mathcal{H}^{2}\left(1+w\right)}\partial_{k}D^{ki}h^{(1)^{\prime}}\partial_{k}D^{k}_{j}h^{(1)^{\prime}}$
$-2\left[\partial^{i}w^{(1)}\mathcal{H}\partial_{j}\phi^{(1)}+\partial^{i}w^{(1)}\partial_{j}\psi^{(1)^{\prime}}+\frac{1}{4}\partial^{i}w^{(1)}\partial_{k}D^{k}_{j}h^{(1)^{\prime}}\right]$
$+\left(4\psi^{(1)}\delta^{ik}-2D^{ik}h^{(1)}\right)\left[\left(\partial_{j}\partial_{k}-\frac{1}{3}\nabla^{2}\delta_{jk}\right)\left(\phi^{(1)}-\psi^{(1)}\right.\right.$
$\left.\left.+2\mathcal{H}w^{(1)}+w^{(1)^{\prime}}\right)\right].$ (117)